### v1

# Modeling, Inference, and Prediction in Mobility-Based Compartmental Models for Epidemiology

Abstract
Classical compartmental models in epidemiology often struggle to accurately capture real-world dynamics due to their inability to address the inherent heterogeneity of populations. In this paper, we introduce a novel approach that incorporates heterogeneity through a mobility variable, transforming the traditional ODE system into a system of integro-differential equations that describe the dynamics of population densities across different compartments. Our results show that, for the same basic reproduction number, our mobility-based model predicts a smaller final pandemic size compared to classic compartmental models, whose population densities are represented as Dirac delta functions in our density-based framework. This addresses the overestimation issue common in many classical models. Additionally, we demonstrate that the time series of the infected population is sufficient to uniquely identify the mobility distribution. We reconstruct this distribution using a machine-learning-based framework, providing both theoretical and algorithmic support to effectively constrain the mobility-based model with real-world data.
1 Introduction
Understanding and predicting epidemic dynamics are crucial for public health planning and response [14, 9, 16]. Accurate models facilitate the timely implementation of control measures, resource allocation, and strategy development, which are essential for minimizing the impact of infectious diseases. For instance, during the COVID-19 pandemic, models by Ferguson et al. [7] guided decisions on lockdowns and social distancing, significantly influencing policy and saving lives.
Compartmental models have been a cornerstone in epidemiology since the early 20th century. These models, such as the classic SIR (Susceptible-Infected-Recovered) model introduced by Kermack and McKendrick [11], divide the population into compartments based on disease status. The transitions between compartments are governed by differential equations that describe the dynamics of disease transmission and recovery. These models have been instrumental in understanding various infectious diseases, including influenza, measles, and more recently, COVID-19. Compartmental models provide a simplified yet powerful framework to study the spread of diseases within a population.
Traditional compartmental models, despite their utility, face significant limitations [10, 1]. One major constraint lies in their assumption of a homogeneous mixing population, where every individual has an equal probability of coming into contact with any other individual, disregarding the complex realities of social interactions and geographic distributions [2, 15]. Furthermore, these models typically omit spatial dynamics, which is critical for understanding disease spread across different regions and the localized impact of interventions [6]. Additionally, traditional models often fail to account for dynamic changes in human behavior during epidemics, such as increased handwashing, social distancing, and vaccination uptake, which can profoundly influence epidemic outcomes [5, 8]. This oversimplification neglects the diverse patterns of human contact within communities.
The current ODE-based epidemiological models frequently overestimate pandemic size by assuming homogeneous mixing and neglecting real-world contact patterns [12, 13, 3]. Moreover, the classic models often fail to incorporate behavioral changes such as increased hygiene practices or timely interventions, which can inflate estimates of susceptible populations and contribute to the overall overestimation. To address these issues, we propose mobility-based compartmental models where the population exhibits heterogeneous mobility distributions. Individuals with higher mobility are at increased risk of infection and transmission during the early stages of the pandemic, potentially leading to a higher basic reproduction number ($\mathcal{R}_{0}$) initially. However, as the pandemic progresses, the initially infected high-mobility group recovers and subsequently provides indirect protection to the low-mobility group, thereby reducing $\mathcal{R}_{0}$ and ultimately limiting the pandemic size. This mobility-based infection mechanism offers a straightforward yet effective approach to more accurately estimate pandemic spread, addressing the overestimation observed in traditional compartmental models used in epidemiology.
Our paper proceeds as follows. In section 2, we propose a mobility-based heterogeneous compartmental model, highlighting its distinctions from homogeneous models in terms of $\mathcal{R}_{0}$ and the final pandemic size. In section 3, we present an inverse problem aimed at inferring the mobility distribution from the time series of the infected population. We provide sufficient conditions to infer the mobility distribution uniquely and propose a machine-learning approach for inferring the mobility distribution. We conclude in section 4.
2 Model description and properties
2.1 A mobility-based SIRS model
The baseline SIRS model delineates infectious disease dynamics within a population, categorizing individuals into susceptible (S), infected (I), and recovered (R) compartments. Consider a system with $N$ agents, in which agent $i$ holds a state $X_{i}$ that takes a value S, I, or R. We define $S(t)$, $I(t)$, and $R(t)$ as the ratios of agents in the three states as
$S(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{S}\}},\quad I(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{I}\}},\quad R(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{R}\}}.$ (1)
The baseline SIRS model posits that individuals transition cyclically between S, I, and R states through infection, recovery, and subsequent loss of immunity, all modeled as independent Poisson-point processes with constant rates. The well-known SIRS model describes the population dynamics of the three compartments as
$\dot{S}(t)$ $=-{\beta}S(t)I(t)+\varepsilon R(t)$ (2)
$\dot{I}(t)$ $={\beta}S(t)I(t)-\gamma I(t)$
$\dot{R}(t)$ $=\gamma I(t)-\varepsilon R(t),$
where $\beta$ is the infection rate, $\gamma$ is the recovery rate, and $\varepsilon$ is the loss of immunity rate.
However, this baseline model overlooks the heterogeneity among individuals within a population, which is critical for understanding disease transmission. Among various factors, mobility plays a pivotal role in epidemics by accelerating infection movement across regions through travel, commuting, and migration. We propose a mobility-based SIRS model where each agent $i$ has a time-independent mobility $m_{i}\in(0,1)$ that determines the infection and transmission rates of agent $i$. All S-I pairs (e.g., a pair of agents $i$ and $j$ with $X_{i}=\text{S}$ and $X_{j}=\text{I}$) are subject to internal infections, with the probability of agent $j$ infecting agent $i$ given by $\beta m_{i}m_{j}$. Infected agents recover at a constant rate $\gamma$, and recovered agents lose immunity at a constant rate $\varepsilon$, consistent with the baseline model (2). Figure 1 illustrates the transitions between the three compartments and their respective rates.
Figure 1: A schematic diagram to show the transition between three compartments and the transition rates in the mobility-based compartments models.
We describe the dynamics using three population density functions of mobility—$S(m,t)$, $I(m,t)$, $R(m,t)$—associated with the susceptible, infected, and recovered compartments, respectively. Here, $S(m,t)\mathrm{d}m,$ represents the ratio of susceptible agents with mobility in the interval $[m,m+\mathrm{d}m,)$. Due to population conservation, we have
$S(m,t)+I(m,t)+R(m,t)=f(m),\quad\text{for all}\leavevmode\nobreak\ t\geq 0,$ (3)
where $f(m)$ denotes the time-independent mobility distribution function. The corresponding mobility-based SIRS model is
$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},+\varepsilon R(m,t)$ (4a)
$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},-\gamma I(m,t)$ (4b)
$\frac{\partial}{\partial t}R(m,t)$ $=\gamma I(m,t)-\varepsilon R(m,t).$ (4c)
In the mobility-based SIRS model, a susceptible agent with mobility $m$ can be infected by any infected agent with mobility $m^{\prime}$, leading to an integral representation of the internal infection process. In our model, each agent’s mobility is time-independent, resulting in local terms for the recovery and loss of immunity processes.
If all agents have the same mobility $m_{0}$, then the mobility distribution is a Dirac delta function centered at $m_{0}$ (i.e., $f(m)=\delta(m-m_{0})$). The mobility-based SIRS model (4) becomes the baseline SIRS model (2) with the infection rate $\beta_{\text{eff}}=\beta m_{0}^{2}$, where we refer $\beta_{\text{eff}}$ to as the effective infection rate.
2.2 Dynamics of the mobility-based SIRS model
In Figure 2, we illustrate the time evolution of the three population densities in the mobility-based SIRS model (4) during the first infection wave. Initially, the population is evenly distributed across the mobility domain $[0,1]$. As the simulation begins, susceptible agents with high mobility are infected first, leading to a peak in the infected population density near $m=1$ at $t=250$. Over time, the infected population recovers, resulting in an increased density in $R(m,t)$. This infection-recovery process occurs earlier in the high-mobility region compared to the low-mobility region.
(a) $S(m,t)$
(b) $I(m,t)$
(c) $R(m,t)$
Figure 2: Evolution dynamics of the mobility-based SIRS model (3), with parameter $\beta=0.1,\varepsilon=0.01$, $\gamma=0.3$ and initial $S(m,0)=0.99,I(m,0)=0.01,R(m,0)=0$, for (a) the susceptible population $S(m,t)$, (b) the infected population $I(m,t)$, and (c) the recovered population $R(m,t)$.
The initial mobility distribution plays a crucial role in determining the mobility heterogeneity of the population as well as shaping the aggregated dynamics. We define the short-hand notation for the integral
$\langle g\rangle(t):=\int_{0}^{1}g(m,t)\leavevmode\nobreak\ \mathrm{d}m,$ (5)
and illustrate the time evolution of the ratios of three compartments in Figure 3. We fix the second-order moment of the initial mobility distribution (i.e., $\langle m^{2}f\rangle$) in all simulations, which, as we will show in section 2.3, is proportional to the basic reproduction number $\mathcal{R}_{0}$. In addition, we use the initial conditions
$S(m,0)=(1-I_{0})f(m),\quad I(m,0)=I_{0}f(m),\quad R(m,0)=0,$ (6)
which we refer to as the proportional initial condition. We will use this initial condition by default without further explanation. We choose three initial mobility distributions that all have the same second-order moment $\langle m^{2}f\rangle$, where $f_{1}$ and $f_{2}$ are restricted normal distributions on $[0,1]$ with
$f_{1}\propto 0.8\mathcal{N}(0.1,0.02)+0.2\mathcal{N}(0.95,0.02),\quad f_{2}\propto\mathcal{N}(0.2,0.365),$ (7)
and $f_{3}$ is a Direct delta function $f(m)=\delta(m-0.435)$.
(a) $\langle S\rangle(t)$
(b) $\langle I\rangle(t)$
(c) $\langle R\rangle(t)$
[…]
2.3 Basic reproduction number
The basic reproduction number $\mathcal{R}_{0}$ is a key epidemiological metric that represents the average number of secondary infections produced by a single infected individual in a fully susceptible population. In this section, we derive $\mathcal{R}_{0}$ of the mobility-based SIRS model (4).
We consider the next-generation matrix method [4] and discretize $I(m,t)$ and $S(m,t)$ on the grids $\{m_{i}\}_{i=1}^{M}$. We denote the discretized quantities by two vectors $\bm{x},\bm{y}\in\mathbb{R}^{M}$, in which
$x_{i}(t)=I(m_{i},t),\quad y_{i}(t)=S(m_{i},t).$ (8)
We consider a numerical quadrature to replace the integral in (4) with
$\int_{0}^{1}m^{\prime}I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},\approx\sum_{i=1}^{M}w_{i}m_{i}x_{i},$ (9)
where $w_{i}$ is the weight of the numerical quadrature at the grid point $m_{i}$. We write (4) in a discrete format and obtain
$\dot{\bm{x}}=\bm{F}(\bm{x},\bm{y})-\bm{V}(\bm{x},\bm{y}),$ (10)
where $\bm{F}$ is the rate of new infections from the susceptible population and $\bm{V}$ is the rate of loss of infections due to recovery. They are
$F_{i}=\beta m_{i}y_{i}\sum_{j}w_{j}m_{j}x_{j},\quad V_{i}=\gamma x_{i}.$ (11)
Consider the disease-free equilibrium, in which the whole population only contains susceptible individuals and its population density satisfies $S(m,t)\equiv f(m)$. Therefore, the disease-free equilibrium vectors $\bm{x}^{0},\bm{y}^{0}\in\mathbb{R}^{M}$ satisfy $x^{0}_{i}=f(m_{i})$ and $y^{0}_{i}=0$. We compute the Jacobi matrices of $\bm{F}$ and $\bm{V}$ with respect to $\bm{x}$ at the disease-free equilibrium vectors $\bm{x}^{0},\bm{y}^{0}$ and obtain
$\frac{\partial\bm{F}}{\partial\bm{x}}(\bm{x}^{0},\bm{y}^{0})_{ij}=\beta m_{i}y^{0}_{i}m_{j}w_{j},\quad\frac{\partial\bm{V}}{\partial\bm{x}}(\bm{x}^{0},\bm{y}^{0})_{ij}=\gamma\delta_{ij},$ (12)
where $\delta_{ij}$ is the Kronecker delta function. We further compute the next-generation matrix
$\frac{\partial\bm{F}}{\partial\bm{x}}\left(\frac{\partial\bm{V}}{\partial\bm{x}}\right)^{-1}=\frac{\beta}{\gamma}\begin{pmatrix}m_{1}y^{0}_{1}\\ \vdots\\ m_{M}y^{0}_{M}\end{pmatrix}\begin{pmatrix}m_{1}w_{1}&\cdots&m_{M}w_{M}\end{pmatrix}_{j}.$ (13)
We compute the largest eigenvalue of the next-generation matrix and obtain the basic reproduction number $\mathcal{R}_{0}=\beta\sum_{i}w_{i}m_{i}^{2}y^{0}_{i}/\gamma$. Notice that $w_{i}$ are the weights of grid points in the numerical quadrature, so as the number of grids $M$ goes to infinite, $\mathcal{R}_{0}$ converges to the integral
$\mathcal{R}_{0}=\frac{\beta}{\gamma}\int_{0}^{1}m^{2}f(m)\leavevmode\nobreak\ \mathrm{d}m,.$ (14)
The above derivations can be made rigorous by using the Fréchet derivative with an appropriate definition of the function space. When the entire population has a constant mobility $m_{0}$, equivalent to the baseline SIRS model (2), the basic reproduction number in (14) is
$\mathcal{R}_{0}=\frac{\beta m_{0}^{2}}{\gamma}=\frac{\beta_{\text{eff}}}{\gamma},$ (15)
which is consistent with the results of the baseline SIRS model (2).
Let $\text{var}(f)$ denote the variance of a mobility distribution $f$. Assuming that all distributions that have the same $\mathcal{R}_{0}$, from (14) we have
$\langle mf\rangle^{2}=\frac{\gamma}{\beta}\mathcal{R}_{0}-\text{var}(f)\leq\frac{\gamma}{\beta}\mathcal{R}_{0}=\langle mf_{\delta}\rangle^{2}.$ (16)
where $f_{\delta}$ is the Dirac delta distribution that has the same second-order moment as other distributions. Therefore, if we fix $\mathcal{R}_{0}$, the Dirac delta distribution $f_{\delta}$ maximizes the mean mobility $\langle mf\rangle$. In practice, we often estimate $\mathcal{R}_{0}$ by considering the early growth rate of cases and performing contact tracing studies, which is independent of model selection. As a result, using the value of $\mathcal{R}_{0}$ to estimate the pandemic size with the baseline SIRS model tends to result in overestimation.
2.4 Mean mobility
Mobility plays a crucial role in the mobility-based compartmental models. We examine the mean mobility of three populations. We focus on the mobility-based SIR model, which is defined by the following equations
$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},$ (17a)
$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},-\gamma I(m,t)$ (17b)
$\frac{\partial}{\partial t}R(m,t)$ $=\gamma I(m,t).$ (17c)
We define $Q_{k}^{\varphi}$ as the $k$th moment with respect to a nonnegative function $\varphi(m,t)$, which is
$Q_{k}^{\varphi}(t)=\int_{0}^{1}m^{k}\varphi(m,t)\leavevmode\nobreak\ \mathrm{d}m.$ (18)
Theorem 1.
Let ${Q_{k}^{S}}(t)$ be the $k$th moment of $S(m,t)$ which is the solution of the mobility-based SIR model (17). If $f$ is not a Dirac delta distribution, then for any $k\geq 0$, the ratio ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases with respect to $t$.
Proof.
We multiply $m^{k}$ to (17a) and integrate $m$ over $[0,1]$. We obtain
$\dot{Q}_{k}^{S}=-\beta Q_{1}^{I}Q_{k+1}^{S},\quad\text{for all}\leavevmode\nobreak\ k\geq 0.$ (19)
By using (19), we obtain the time derivative of ${Q_{k+1}^{S}}/{Q_{k}^{S}}$
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{Q_{k+1}^{S}}{Q_{k}^{S}}\right)=\frac{1}{Q_{k}^{S}Q_{k}^{S}}\left(Q_{k}^{S}\dot{Q}_{k+1}^{S}-\dot{Q}_{k}^{S}Q_{k+1}^{S}\right)=-\frac{\beta Q_{1}^{I}}{Q_{k}^{S}Q_{k}^{S}}\left(Q_{k}^{S}Q_{k+2}^{S}-Q_{k+1}^{S}Q_{k+1}^{S}\right).$ (20)
From Hölder’s inequality, we know that
$Q_{k}^{S}Q_{k+2}^{S}-Q_{k+1}^{S}Q_{k+1}^{S}>0,$ (21)
which implies that the time derivative in (20) is negative, so that ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases. ∎
For $k=0$, the ratio in Theorem 1 simplifies to $Q_{1}^{S}(t)/Q_{0}^{S}(t)$, representing the mean mobility of the susceptible population. Thus, in the mobility-based SIR model (17), the mean mobility of the susceptible population decreases over time $t$. However, this monotonic property does not necessarily apply to the infected and recovered populations, as illustrated in Figure 4. In addition, if $f$ is a Dirac delta distribution, then the mean mobility remains constant and is independent of time $t$.
(a) $Q_{1}^{S}/Q_{0}^{S}$
(b) $Q_{1}^{I}/Q_{0}^{I}$
(c) $Q_{1}^{R}/Q_{0}^{R}$
[…]
2.5 Final pandemic size
The final pandemic size refers to the total number of individuals infected throughout an epidemic and is a crucial quantity to assess disease transmission dynamics. For the classical SIR model with a basic reproduction number $\mathcal{R}_{0}$, it is well known that the final pandemic size $R_{\infty}$ satisfies the equation
$R_{\infty}+\exp(-\mathcal{R}_{0}R_{\infty})=1,$ (22)
if the initial infected population $I\approx 0$. In the real world, the population exhibits heterogeneous mobility, with high-mobility individuals likely being infected first during the early pandemic stages. Once recovered, they provide some protection to low-mobility individuals. In practice, $\mathcal{R}_{0}$ is typically estimated from the initial exponential growth rate, which is biased toward the infection rate of high-mobility individuals, thereby resulting in an overestimation of the final pandemic size. In this section, we derive the final pandemic size of the mobility-based SIR model (17) with the proportional initial condition (6). We rigorously demonstrate why the classic SIR model tends to overestimate the final size of pandemics, assuming the same $\mathcal{R}_{0}$ as in the mobility-based SIR model.
Theorem 2.
Let $R_{\infty}$ be the final pandemic size of the mobility-based SIR model (17) with the proportial initial condition (6). Assume $I=0\approx 0$, then $R_{\infty}$ satisfies
$R_{\infty}+\langle f\exp\left(-\alpha m\beta/\gamma\right)\rangle$ $=1,$ (23)
where $\alpha$ is the solution of $\alpha+\langle mf\exp\left(-\alpha m\beta/\gamma\right)\rangle=\langle mf\rangle$ on the interval $[0,1]$.
Proof.
By solving (2) for $S(m,t)$, we obtain that
$S(m,t)=S(m,0)\exp\left(-\frac{\beta}{\gamma}m\langle mR\rangle(t)\right).$ (24)
We multiply the above equation by $m$ and integrate it over $m$, then obtain
$\langle mS\rangle(t)=\langle mS(m,0)\exp\left(-\frac{\beta}{\gamma}m\langle mR\rangle(t)\right)\rangle.$ (25)
Since $I(m,t)\to 0$ as $t\to\infty$, we have
$\langle mf\rangle-\langle mR\rangle(\infty)=\langle mS(m,0)\exp\left(-\frac{\beta}{\gamma}m\langle mR\rangle(\infty)\right)\rangle.$ (26)
We use the initial condition (6) and assume that $S(m,0)\approx f(m)$, then $\alpha=\langle mR\rangle(\infty)$ satisfies
$\alpha+\langle mf\exp\left(-\alpha m\beta/\gamma\right)\rangle=\langle mf\rangle.$ (27)
Integrating (24) and taking $t$ to infinite, we obtain that $R_{\infty}=\langle R\rangle(\infty)$ satisfies (23). ∎
Theorem 2 offers a method to compute the final pandemic size without the need to simulate the model (17) over time. The process involves first determining $\alpha$ by finding the root of (27) within the interval $[0,1]$. Once $\alpha$ is obtained, we then solve (23) to find $R_{\infty}$. Additionally, once the parameters $\beta$ and $\gamma$ are fixed, the mobility distribution $f(m)$ solely determines both the final pandemic size $R_{\infty}$ and the basic reproduction number $\mathcal{R}_{0}$ as given in (14). We will demonstrate in Theorem 3 that among all distributions that yield the same basic reproduction number, the Dirac delta distribution $f_{\delta}$, which corresponds to a homogeneous mobility of the population, maximizes the final pandemic size.
Theorem 3.
Let $R_{\infty}$ and $R^{\delta}_{\infty}$ be the final pandemic sizes associated with a distribution $f$ and a Dirac delta distribution $f_{\delta}$, respectively. Suppose that $f$ and $f_{\delta}$ yield the same basic reproduction number $\mathcal{R}_{0}$ (14), it holds that $R_{\infty}\leq R^{\delta}_{\infty}$ with the equality occurring only if $f=f_{\delta}$.
Theorem 2 provides an approach to compute the final pandemic size. We introduce another perspective to represent the final pandemic size to prove Theorem 3.
Proof.
From (17), we have
$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}t}=-\beta Q_{1}^{S}Q_{1}^{I},\quad\frac{\mathrm{d}Q_{0}^{R}}{\mathrm{d}t}=\gamma Q_{0}^{I},\quad t\in[0,\infty).$ (28)
Since $Q_{0}^{R}(t)$ monotonely increases with respect to $t$, so we have a one-to-one correspondence between $Q_{0}^{R}(t)$ and $t$. Therefore, we can view all functions of $t$ (including $Q_{0}^{S}$) as a function of $Q_{0}^{R}$. Equation (28) becomes
$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}Q_{0}^{R}}=-\frac{\beta}{\gamma}\widetilde{H}(Q_{0}^{R})Q_{0}^{S},\quad Q_{0}^{R}\in[Q_{0}^{R}(0),Q_{0}^{R}(\infty)),$ (29)
where
$\widetilde{H}(Q_{0}^{R})=H(t)=\frac{Q_{1}^{S}(t)Q_{1}^{I}(t)}{Q_{0}^{S}(t)Q_{0}^{I}(t)}.$ (30)
Recall that $Q_{0}^{R}(0)=0$. We integrate (29) over the interval $[0,Q_{0}^{R}(\infty)]$ and obtain
$Q_{0}^{S}(\infty)=\exp\left(-\frac{\beta}{\gamma}\int_{0}^{Q_{0}^{R}(\infty)}\widetilde{H}(\alpha)\leavevmode\nobreak\ \mathrm{d}\alpha\right).$ (31)
Using $\lim_{t\to\infty}Q_{0}^{I}(t)=0$ and change of variable, we obtain
$Q_{0}^{R}(\infty)+\exp\left(-\frac{\beta}{\gamma}\int_{0}^{\infty}H(t)\leavevmode\nobreak\ \mathrm{d}Q_{0}^{R}(t)\right)=1.$ (32)
As $R_{\infty}^{\delta}$ satisfies
$R^{\delta}_{\infty}+\exp\left(-\mathcal{R}_{0}R^{\delta}_{\infty}\right)=1,$ (33)
it is sufficient to prove that
$H(t)\leq\frac{\gamma}{\beta}\mathcal{R}_{0}=\langle m^{2}f\rangle,\quad\text{for all\leavevmode\nobreak\ }t>0,$ (34)
where the equal sign holds only if $f=f_{\delta}$. It is not difficult to verify that the equality holds when $f=f_{\delta}$. We will assume that $f\neq f_{\delta}$ and prove that $H(t)<\langle m^{2}f\rangle$ in the following proof.
From Theorem 1 for $k=1$ and Lemma 4, we have
$\frac{Q_{1}^{S}(t)}{Q_{0}^{S}(t)}<\langle mf\rangle\quad\text{and}\quad\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{\langle m^{2}f\rangle}{\langle mf\rangle},\quad\text{for all\leavevmode\nobreak\ }t>0.$ (35)
Combining it with (30), we complete the proof. ∎
Lemma 4.
Let $I(m,t)$ be the solution of the mobility-based SIR model (17) with the proportional initial condition (6). For all $t>0$, it holds that
$\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{\langle m^{2}f\rangle}{\langle mf\rangle},$ (36)
with equality occurring only if $f$ is a Dirac delta distribution.
Proof.
If $f$ is a Dirac delta distribution $f(m)=\delta(m-m_{0})$, then both sides of (36) equal $m_{0}$. We assume that $f$ is not a Dirac delta distribution in the following proof.
Using (17), we compute the time derivative of ${Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}$ and obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{Q_{1}^{I}}{Q_{0}^{I}}\right)=\frac{\beta Q_{1}^{I}Q_{1}^{S}}{Q_{0}^{I}}\left(\frac{Q_{2}^{S}}{Q_{1}^{S}}-\frac{Q_{1}^{I}}{Q_{0}^{I}}\right).$ (37)
We denote the instants that $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}\right)$ changes its sign as $t_{k}$ and partition the time interval $(0,\infty)$ into open intervals $\Omega_{k}$ with $t_{k}$ as indicated in Figure 5.
Figure 5: A partition of intervals due to the sign of (37).
Using the initial condition (6), we have
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{Q_{1}^{I}}{Q_{0}^{I}}\right)(0)=\frac{\beta Q_{1}^{I}(0)Q_{1}^{S}(0)}{Q_{0}^{I}(0)}\left(\frac{\langle m^{2}f\rangle}{\langle mf\rangle}-\langle mf\rangle\right)>0,$ (38)
which implies that $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}}/{Q_{0}^{I}}\right)$ is nonnegative on the interval $\Omega_{1}$. Due to the smoothness, the signs of (37) interlace on $\Omega_{1}$, $\Omega_{2}$, $\Omega_{3}$, $\cdots$.
If $t\in\cup_{k=1}^{\infty}\overline{\Omega_{2k-1}}$, then we know (37) is nonnegative, which implies that
$\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{Q_{2}^{S}(t)}{Q_{1}^{S}(t)}<\frac{Q_{2}^{S}(0)}{Q_{1}^{S}(0)}=\frac{\langle m^{2}f\rangle}{\langle mf\rangle},$ (39)
where the second inequality uses Theorem 1 for the case $k=1$. If $t\in\cup_{k=1}^{\infty}\Omega_{2k}$ (supposing $t\in\Omega_{2k}$), then ${Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}$ decreases with respect to $t$ on the interval $\Omega_{2k}$. Due to continuity, we have
$\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{Q_{1}^{I}(t_{2k-1})}{Q_{0}^{I}(t_{2k-1})}<\frac{\langle m^{2}f\rangle}{\langle mf\rangle},\quad\text{for all\leavevmode\nobreak\ }t\in\Omega_{2k-1},$ (40)
where the second inequality holds from (39).
3 Learning the mobility distribution from data
Mobility holds critical significance in epidemic models as it directly influences the spread and dynamics of infectious diseases within populations. Understanding how people move between locations allows for more accurate predictions of disease transmission and enhances the precision of epidemic models to reflect real-world scenarios. However, the mobility of a population involves complex factors, including travel patterns, migration, and commuting, which are often impossible to measure directly in real-world situations. Realistically, the available data in epidemics is the number of infected individuals, which translates to the ratio of infected agents at time $t$:
$Q_{0}^{I}(t)=\int_{0}^{1}I(m,t)\leavevmode\nobreak\ \mathrm{d}m,$ (41)
in the mobility-based SIRS model (4). In this section, we infer the mobility distribution $f(m)\in\mathcal{L}^{1}(0,1)$ from the time series of $Q_{0}^{I}(t)$, offering both a theoretical guarantee of unique inference and a machine-learning-based numerical approach.
3.1 Theoretical guarantees
Theorem 5.
Let $I(m,t)$ be the infected mobility distribution of the mobility-based SIRS model (4) with the initial condition (6), yielding the 0th-order moment $Q_{0}^{I}(t)$. Assume that $Q_{0}^{I}(m,t)$ is arbitrarily smooth (i.e., $Q_{0}^{I}\in\mathcal{C}^{\infty}(0,\infty)$) and all the parameters ($I_{0}$, $\beta$, $\gamma$, and $\varepsilon$) are fixed and not equal to 0, then there is a one-to-one correspondence between $f$ and $Q_{0}^{I}$.
This theorem indicates that for any two different mobility distributions ($f\neq\tilde{f}$), the corresponding time series of the infected numbers ($Q_{0}^{I}$ and $Q_{0}^{\tilde{I}}$) also differ. This implies that it is sufficient to infer $f$ from $I$ theoretically and this inference problem has a unique solution. We prove this theorem at the end of this subsection.
Recall the moment notation in (18). We multiply $m^{k}$ to (4) and obtain the equations for moments
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{S}$ $=-\beta Q_{1}^{I}Q_{k+1}^{S}+\varepsilon Q_{k}^{R}$ (42a)
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{I}$ $=\beta Q_{1}^{I}Q_{k+1}^{S}-\gamma Q_{k}^{I}$ (42b)
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{R}$ $=\gamma Q_{k}^{I}-\varepsilon Q_{k}^{R}.$ (42c)
Consider another set of solutions $\tilde{S}(m,t),\tilde{I}(m,t),\tilde{R}(m,t)$ of (4) with the initial condition proportional to $\tilde{f}(m)$.
Lemma 6.
The following three statements are equivalent:
$Q_{k}^{S}=Q_{k}^{\tilde{S}},\quad Q_{k}^{I}=Q_{k}^{\tilde{I}},\quad\text{and}\quad Q_{k}^{R}=Q_{k}^{\tilde{R}}.$ (43)
Proof.
We add (42a) and (42b) and obtain
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{S}+\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{I}=\varepsilon Q_{k}^{f}-(\varepsilon+\gamma)Q_{k}^{I}-\varepsilon Q_{k}^{S},$ (44)
which utilizes the fact that $Q_{k}^{f}=Q_{k}^{S}+Q_{k}^{I}+Q_{k}^{R}$. Similarly, we have
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{\tilde{S}}+\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{\tilde{I}}=\varepsilon Q_{k}^{\tilde{f}}-(\varepsilon+\gamma)Q_{k}^{\tilde{I}}-\varepsilon Q_{k}^{\tilde{S}}.$ (45)
Step 1. We first show that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$ is equivalent to $Q_{k}^{S}=Q_{k}^{\tilde{S}}$. Suppose that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$. We compute the difference between (44) and (45) and obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\left(Q_{k}^{S}-Q_{k}^{\tilde{S}}\right)=\varepsilon(Q_{k}^{f}-Q_{k}^{\tilde{f}})-\varepsilon(Q_{k}^{S}-Q_{k}^{\tilde{S}}).$ (46)
Since $Q_{k}^{I}(0)=I_{0}Q_{k}^{f}$ and $Q_{k}^{\tilde{I}}(0)=I_{0}Q_{k}^{\tilde{f}}$, we use $Q_{k}^{I}(0)=Q_{k}^{\tilde{I}}(0)$ and obtain
$Q_{k}^{f}=Q_{k}^{\tilde{f}},\quad Q_{k}^{S}(0)-Q_{k}^{\tilde{S}}(0)=S_{0}\left(Q_{k}^{f}-Q_{k}^{\tilde{f}}\right)=0.$ (47)
Therefore, $e(t)=Q_{k}^{S}(t)-Q_{k}^{\tilde{S}}(t)$ satisfies a linear ODE
$\frac{\mathrm{d}}{\mathrm{d}t}e(t)=-\varepsilon e(t),\quad e(0)=0,$ (48)
yielding $e(t)=0$ for all $t\geq 0$, so that $Q_{k}^{S}=Q_{k}^{\tilde{S}}$. Suppose that $Q_{k}^{S}=Q_{k}^{\tilde{S}}$. We use the similar idea as above and show $e(t)=Q_{k}^{I}(t)-Q_{k}^{\tilde{I}}(t)$ satisfies
$\frac{\mathrm{d}}{\mathrm{d}t}e(t)=-(\varepsilon+\gamma)e(t),\quad e(0)=0,$ (49)
which yields $Q_{k}^{I}(t)=Q_{k}^{\tilde{I}}(t)$ for all $t\geq 0$.
Step 2. We then show that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$ is equivalent to $Q_{k}^{R}=Q_{k}^{\tilde{R}}$. Suppose that $Q_{k}^{R}=Q_{k}^{\tilde{R}}$. Therefore, we have $\mathrm{d}Q_{k}^{R}/\mathrm{d}t=\mathrm{d}Q_{k}^{\tilde{R}}/\mathrm{d}t$. From (42c), we have $Q_{k}^{I}=Q_{k}^{\tilde{I}}$. Suppose $Q_{k}^{I}=Q_{k}^{\tilde{I}}$. Using the similar idea as above, we have $e(t)=Q_{k}^{R}-Q_{k}^{\tilde{R}}$ satisfies
$\frac{\mathrm{d}}{\mathrm{d}t}e(t)=-\varepsilon e(t),\quad e(0)=0,$ (50)
implying that $Q_{k}^{R}=Q_{k}^{\tilde{R}}$. ∎
Theorem 5 and Lemma 6 indicate that there is a one-to-one correspondence between $f$ and $Q_{0}^{S}$ and that between $f$ and $Q_{0}^{R}$. This implies that, alternatively, we can also infer the mobility distribution $f$ from the time series of $Q_{0}^{S}$ or $Q_{0}^{R}$ if they are available and the inference problems yield unique solutions.
Lemma 7.
Let $L$ be a positive integer and the moments satisfy
$\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{S,I,R}(0)$ $=\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad k=0,1,\ldots,K-1-\ell,$ (51)
$\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{S,I,R}(0)$ $\neq\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad k=K-\ell.$
for $\ell=0,1,\ldots,L-1$, then (51) also holds for $\ell=L$.
Proof.
A direct computation yields
$\frac{\mathrm{d}^{L}}{\mathrm{d}t^{L}}Q_{k}^{I}=\frac{\mathrm{d}^{L-1}}{\mathrm{d}t^{L-1}}(\beta Q_{1}^{I}Q_{k+1}^{S}-\gamma Q_{k}^{I})=\beta\sum_{\alpha=0}^{{L-1}}C_{L-1}^{\alpha}\frac{\mathrm{d}^{\alpha}}{\mathrm{d}t^{\alpha}}Q_{1}^{I}\frac{\mathrm{d}^{{L-1}-\alpha}}{\mathrm{d}t^{{L-1}-\alpha}}Q_{k+1}^{S}-\gamma\frac{\mathrm{d}^{L-1}}{\mathrm{d}t^{L-1}}Q_{k}^{I}.$ (52)
For $k=0,\ldots,K-L-1$, we use conditions in (51) and obtain
$\frac{\mathrm{d}^{L}}{\mathrm{d}t^{L}}Q_{k}^{I}(0)=\frac{\mathrm{d}^{L}}{\mathrm{d}t^{L}}Q_{k}^{\tilde{I}}(0).$ (53)
For $k=K-L$, we have
$\frac{\mathrm{d}^{L}}{\mathrm{d}t^{L}}Q_{K-L}^{I}$ $=\beta\sum_{\alpha=0}^{{L-1}}C_{L-1}^{\alpha}\frac{\mathrm{d}^{\alpha}}{\mathrm{d}t^{\alpha}}Q_{1}^{I}\frac{\mathrm{d}^{{L-1}-\alpha}}{\mathrm{d}t^{{L-1}-\alpha}}Q_{{K-L}+1}^{S}-\gamma\frac{\mathrm{d}^{L-1}}{\mathrm{d}t^{L-1}}Q_{K-L}^{I}$ (54)
$=\beta\sum_{\alpha=1}^{{L-1}}C_{L-1}^{\alpha}\frac{\mathrm{d}^{\alpha}}{\mathrm{d}t^{\alpha}}Q_{1}^{I}\frac{\mathrm{d}^{{L-1}-\alpha}}{\mathrm{d}t^{{L-1}-\alpha}}Q_{{K-L}+1}^{S}-\gamma\frac{\mathrm{d}^{L-1}}{\mathrm{d}t^{L-1}}Q_{K-L}^{I}+\beta Q_{1}^{I}{\frac{\mathrm{d}^{L-1}}{\mathrm{d}t^{L-1}}Q_{{K-L}+1}^{S}}.$
All paired moments from $f$ and $\tilde{f}$ are equal to each other at $t=0$ from (42), except for the last term since
${\frac{\mathrm{d}^{L-1}}{\mathrm{d}t^{L-1}}Q_{{K-L}+1}^{S}}(0)\neq{\frac{\mathrm{d}^{L-1}}{\mathrm{d}t^{L-1}}Q_{{K-L}+1}^{\tilde{S}}}(0).$ (55)
We complete the proof for moments for the infected population. We repeat similar calculations for the moments of $S$ and $R$ and obtain the results. ∎
Now we prove Theorem 5.
Proof.
We use the same moment notation (18). Since $f\neq\tilde{f}$, we can find a positive integer $K$ such that
$Q_{k}^{f}=Q_{k}^{\tilde{f}},\quad\text{\leavevmode\nobreak\ for all}\leavevmode\nobreak\ k=0,1,\ldots,K-1,$ (56)
$Q_{K}^{f}\neq Q_{K}^{\tilde{f}}.$
From the proportional initial condition (6), we have
$Q_{k}^{S,I,R}(0)=Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad\text{for all}\leavevmode\nobreak\ k=0,1,\ldots,K-1,$ (57)
$Q_{K}^{S,I,R}(0)\neq Q_{K}^{\tilde{S},\tilde{I},\tilde{R}}(0).$
We plug the initial values into (42) and obtain
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{S,I,R}(0)=\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad\text{for all}\leavevmode\nobreak\ k=0,1,\ldots,K-2,$ (58)
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{K-1}^{S,I,R}(0)\neq\frac{\mathrm{d}}{\mathrm{d}t}Q_{K-1}^{\tilde{S},\tilde{I},\tilde{R}}(0),$
where the last inequality for the recovered population also uses the facts that
$Q_{K-1}^{f}=Q_{K-1}^{S}(0)+Q_{K-1}^{I}(0)+Q_{K-1}^{R}(0),\quad Q_{K-1}^{\tilde{f}}=Q_{K-1}^{\tilde{S}}(0)+Q_{K-1}^{\tilde{I}}(0)+Q_{K-1}^{\tilde{R}}(0).$ (59)
Notice that (58) satisfies the conditions in Lemma 7 for $L=2$. By induction, we know (51) hold for $\ell=2,3,\ldots,K$. In particular, we use the inequality result in Lemma 7 for $\ell=K$ and $k=0$ yielding
$\frac{\mathrm{d}^{K}}{\mathrm{d}t^{K}}Q_{0}^{I}(0)\neq\frac{\mathrm{d}^{K}}{\mathrm{d}t^{K}}Q_{0}^{\tilde{I}}(0).$ (60)
Therefore, we have $Q_{0}^{I}\neq Q_{0}^{\tilde{I}}$ and the correspondence between $f$ and $Q_{0}^{I}$ is one-to-one.
3.2 Learning the mobility distribution using a neural-network framework
Mobility distribution significantly impacts the pandemic size and transmission. However, in reality, it is not feasible to directly measure the mobility density functions of different compartments during pandemics. Realistically, the available data in epidemics is often limited to the number of infected individuals. Theorem 5 suggests a one-to-one correspondence between the mobility distribution $f(m)$ and the ratio of the infected population $\langle I\rangle(t)$. Thus, a natural approach is to infer the mobility distribution from the early-stage pandemic data to make accurate predictions for the future course of the pandemic. We propose a neural-network-based approach to learn the operator from the time-series $\langle I\rangle(\cdot)$ to the mobility distribution $f(\cdot)$.
Assume that the pandemic data takes the form of $\mathbf{I}=(\langle I\rangle(t_{1}),\ldots,\langle I\rangle(t_{N_{t}}))^{T}$, which is a time series of the ratio of the infected population at discrete time steps. We discretize $\mathbf{f}$ using the discrete cosine transform (DCF) and denote the first $N_{f}$ modes as $\mathbf{f}=(\hat{f}_{1},\ldots,\hat{f}_{N_{f}})^{T}$. We consider a feedforward neural network $\mathrm{G}_{\theta}$ to learn the mapping between $\mathbf{I}$ and $\mathbf{f}$:
$\mathbf{f}=\mathrm{G}_{\theta}(\mathbf{I}).$ (61)
We sample the mobility distribution $f$ and obtain $\langle I\rangle$ by simulating the mobility-based SIRS model (4) to form the training set. When using neural networks to learn unknown parameters or functions, it is crucial to sample a training set that encompasses a wide range of scenarios. In our context, this involves properly sampling the mobility distribution comprehensively and robustly. Since a Gaussian mixture model is a universal approximator of probability densities, we sample $f$ by a Gaussian mixture with random coefficients. In particular, we let
${f}\propto\sum_{n=1}^{3}r_{n}\mathcal{N}(\mu_{n},\sigma_{n}^{2}),$ (62)
where $\mathcal{N}(\mu,\sigma^{2})$ is the Gaussian density function with mean $\mu$ and variance $\sigma^{2}$ and $r_{n}$, $\mu_{n}$, and $\sigma_{n}$ are all uniformly distributed on $[0,1]$, $[-0.2,1.2]$, and $[0.01,0.26]$, respectively.
Suppose the dataset $\mathcal{D}$ contains $K$ pairs of $\mathbf{I}$ and $\mathbf{f}$, which we obtain by discretizing $\langle I\rangle$ and $f$. We parametrize $\mathrm{G}_{\theta}$ with a fully-connected feedforward neural network and minimize the loss function
$\mathrm{L}(\theta)=\frac{1}{K}\sum_{i=1}^{K}\left[\mathrm{G}_{\theta}(\mathbf{I}_{i})-\mathbf{f}_{i}\right]^{2}.$
In our simulations, we use a six-layer-fully connected neural network. The width of layers are 128, 512, 512, 512, 128, and 32, respectively. Each layer includes a Sigmoid activation function and $L^{2}$ regularization with a magnitude of $0.005$. In addition, we choose $N_{f}=8$ and $N_{t}=101$. For each training step, we sample a random batch from the training set and use the Adam optimizer with a learning rate of 0.0005 to minimize the loss function for each batch. We generate a training set of size 10,000 and train the neural network for 2,000 epochs. During each training epoch, we optimize the loss function using 1,562 batches with a batch size of 54.
3.3 Numerical results
[…]
Given the dataset $\mathcal{D}$, we obtain the mapping $\mathrm{G}_{\theta}$ by minimizing the loss function (3.2). To evaluate performance, we generate an additional six mobility distributions using (62) and simulate the corresponding time series of infected populations $\mathbf{I}$ as inputs. Figure 6 shows a comparison between the learned mobility distributions and the ground truth. It is noteworthy that the testing inputs $\mathbf{I}_{i}$ are not part of the original training set; however, the inferred mobility distributions are capable of capturing the trends of the ground truth and yield time series that closely match the ground truth.
(a) Predicted and true mobility distributions $f(m)$
(b) Predicted and true infected populations $\langle I\rangle(t)$
Figure 6: (a): A comparison between the predicted mobility distributions and the true distributions. (b): A comparison between the predicted and true $\langle I\rangle(t)$.
We analyze mobility distributions using real COVID-19 data, shown in Figure 7. From February 29th, 2020, to September 16th, 2020, covering the initial 200 days of our dataset, we examine the infection ratio between Massachusetts and New York. We calculate the total infected cases by summing daily new infections up to the current time, adjusted with a decaying prefactor to account for previously infected individuals. Specifically, we use
$\langle I\rangle(t)=I_{\text{daily}}(t)+\exp{(-1/\gamma)}\langle I\rangle(t-1).$ (63)
We recover the daily true case counts $I_{\text{daily}}(t)$ from the death data and fit them with exponential growth for the first 30 days. The real data reflects a much smaller scale of disease transmission compared to the training set. To use the real data as input, we scale the input by a factor of 50. We show the resulting input data in Figure 7(a) and display the learned mobility distributions in Figure 7(b).
(a) Real COVID-19 data $\langle I\rangle(t)$
(b) Predicted mobility $f(m)$
Figure 7: (a): The scaled ratio of COVID-19 infected cases during the initial wave in Massachusetts and New York from February 29th, 2020, to September 16th, 2020. (b): Inferred mobility distributions corresponding to the inputs in (a).
4 Conclusion and outlook
In this paper, we integrate social heterogeneity into compartmental models by introducing a mobility variable $m$, assuming it ranges from 00 to $1$ for each individual. This approach represents the population in each compartment (susceptible, infected, recovered, etc.) as a density function over $[0,1]$. Individuals with higher mobility are more prone to transmitting infectious diseases to others. The time evolution of population density for each compartment is described through an integro-differential equation.
We explore the key properties of the proposed mobility-based compartmental models crucial for epidemiological modeling, such as the basic reproduction number $\mathcal{R}_{0}$, final pandemic size, and mean mobility. The proposed mobility-based compartment models also effectively mitigate the overestimation problem inherent in many homogeneous compartmental models when estimating the final pandemic size. The basic reproduction rate $\mathcal{R}_{0}$ is often estimated based on the initial growth rate of the infected population. In the early stages of a pandemic, individuals with higher mobility tend to get infected quickly, contributing to a higher $\mathcal{R}_{0}$. As the pandemic evolves, the recovered population with higher mobility can impart some degree of immunity to the broader population, particularly those with lower mobility. This process helps in mitigating the eventual final size of the pandemic. We rigorously demonstrate that under the same $\mathcal{R}_{0}$, the classical SIR model yields the largest final pandemic size. Numerical findings consistently indicate that a polarized mobility distribution typically results in a significantly smaller final pandemic size compared to the classical SIR model. Additionally, we establish a one-to-one correspondence between the time series of the total infected population and the mobility distribution, demonstrating that the mobility distribution can be practically inferred from infection data. We propose a deep learning-based algorithm for inferring the mobility distribution directly from the total infected population.
The primary objective of integrating mobility into compartmental disease models is to utilize the mobility distribution as an intermediary connecting disease dynamics with social dynamics. Changes in public health policies and shifts in public opinion about infectious diseases influence disease spread by altering people’s mobility patterns. Moving forward, we plan to incorporate opinion dynamics into mobility-based compartmental models to explore the interplay between public opinion and disease transmission dynamics.
Acknowledgement
YL and NJ are partially supported by NSF DMS-2108628.

### v2

@@
 Abstract
-Classical compartmental models in epidemiology often struggle to accurately capture real-world dynamics due to their inability to address the inherent heterogeneity of populations.
-In this paper, we introduce a novel approach that incorporates heterogeneity through a mobility variable, transforming the traditional ODE system into a system of integro-differential equations that describe the dynamics of population densities across different compartments.
-Our results show that, for the same basic reproduction number, our mobility-based model predicts a smaller final pandemic size compared to classic compartmental models, whose population densities are represented as Dirac delta functions in our density-based framework.
-This addresses the overestimation issue common in many classical models.
-Additionally, we demonstrate that the time series of the infected population is sufficient to uniquely identify the mobility distribution.
-We reconstruct this distribution using a machine-learning-based framework, providing both theoretical and algorithmic support to effectively constrain the mobility-based model with real-world data.
+Classical compartmental models in epidemiology often assume a homogeneous population for simplicity, which neglects the inherent heterogeneity among individuals.
+This assumption frequently leads to inaccurate predictions when applied to real-world data.
+For example, evidence has shown that classical models overestimate the final pandemic size in the H1N1-2009 and COVID-19 outbreaks.
+To address this issue, we introduce individual mobility as a key factor in disease transmission and control.
+We characterize disease dynamics using mobility distribution functions for each compartment and propose a mobility-based compartmental model that incorporates population heterogeneity.
+Our results demonstrate that, for the same basic reproduction number, our mobility-based model predicts a smaller final pandemic size compared to the classical models, effectively addressing the common overestimation problem.
+Additionally, we infer mobility distributions from the time series of the infected population.
+We provide sufficient conditions for uniquely identifying the mobility distribution from a dataset and propose a machine-learning-based approach to learn mobility from both synthesized and real-world data.
 1 Introduction
-Understanding and predicting epidemic dynamics are crucial for public health planning and response [14, 9, 16].
-Accurate models facilitate the timely implementation of control measures, resource allocation, and strategy development, which are essential for minimizing the impact of infectious diseases.
-For instance, during the COVID-19 pandemic, models by Ferguson et al.
-[7] guided decisions on lockdowns and social distancing, significantly influencing policy and saving lives.
-Compartmental models have been a cornerstone in epidemiology since the early 20th century.
-These models, such as the classic SIR (Susceptible-Infected-Recovered) model introduced by Kermack and McKendrick [11], divide the population into compartments based on disease status.
-The transitions between compartments are governed by differential equations that describe the dynamics of disease transmission and recovery.
-These models have been instrumental in understanding various infectious diseases, including influenza, measles, and more recently, COVID-19.
-Compartmental models provide a simplified yet powerful framework to study the spread of diseases within a population.
-Traditional compartmental models, despite their utility, face significant limitations [10, 1].
-One major constraint lies in their assumption of a homogeneous mixing population, where every individual has an equal probability of coming into contact with any other individual, disregarding the complex realities of social interactions and geographic distributions [2, 15].
-Furthermore, these models typically omit spatial dynamics, which is critical for understanding disease spread across different regions and the localized impact of interventions [6].
-Additionally, traditional models often fail to account for dynamic changes in human behavior during epidemics, such as increased handwashing, social distancing, and vaccination uptake, which can profoundly influence epidemic outcomes [5, 8].
-This oversimplification neglects the diverse patterns of human contact within communities.
-The current ODE-based epidemiological models frequently overestimate pandemic size by assuming homogeneous mixing and neglecting real-world contact patterns [12, 13, 3].
-Moreover, the classic models often fail to incorporate behavioral changes such as increased hygiene practices or timely interventions, which can inflate estimates of susceptible populations and contribute to the overall overestimation.
-To address these issues, we propose mobility-based compartmental models where the population exhibits heterogeneous mobility distributions.
-Individuals with higher mobility are at increased risk of infection and transmission during the early stages of the pandemic, potentially leading to a higher basic reproduction number ($\mathcal{R}_{0}$) initially.
-However, as the pandemic progresses, the initially infected high-mobility group recovers and subsequently provides indirect protection to the low-mobility group, thereby reducing $\mathcal{R}_{0}$ and ultimately limiting the pandemic size.
-This mobility-based infection mechanism offers a straightforward yet effective approach to more accurately estimate pandemic spread, addressing the overestimation observed in traditional compartmental models used in epidemiology.
+Understanding and predicting epidemic dynamics are essential for effective public health planning and response [23, 14, 26].
+Reliable and effective models provide crucial insights that enable timely implementation of control measures, efficient resource allocation, and strategic planning.
+For example, during the COVID-19 pandemic, models by Ferguson et al.
+[11] and IHME [5] guided decisions on school closures, social distancing, mask mandates, and reopening.
+Epidemiological models significantly influenced policy and played a key role in saving lives.
+Among all these models, compartmental models have served as a cornerstone in epidemiology since the early 20th century.
+These compartmental models, including the well-known SIR (Susceptible-Infected-Recovered) model introduced by Kermack and McKendrick [18] in 1927, divide the individuals into compartments based on their disease status and use ordinary differential equations to describe the dynamic evolution of each compartment.
+Despite their utility, classical compartmental models have significant limitations [17, 1].
+A major issue is their assumption of a homogeneous mixing population, where each individual has an equal chance of interacting with any other.
+This overlooks spatial dynamics [10, 28], real-world contact patterns [20, 21, 3], and dynamic changes in human behavior during epidemics [9, 12], such as increased handwashing, social distancing, and vaccination, which can significantly affect epidemic outcomes.
+The COVID-19 pandemic further highlights that classical models often ignore behavioral changes and interventions, leading to inflated estimates of susceptible populations.
+To address these limitations, we propose mobility-based compartmental models incorporating heterogeneous mobility distributions.
+These models describe the time evolution of population densities for each compartment as a function of a mobility variable ranging from $[0,1]$, making them infinite-dimensional.
+While the literature has explored similar infinite-dimensional models for their basic mathematical properties, to the best of our knowledge, this paper addresses critical topics in epidemiology such as the estimation of basic reproduction number and final pandemic size and the identifiability of the mobility distribution, which have not been previously studied.
+A notable innovation in our paper is the introduction of individual mobility, a key factor in infectious disease spread.
+During an outbreak, individuals with higher mobility face greater risks of infection and transmission early on.
+As the pandemic progresses, this high-risk group recovers and indirectly protects the low-mobility group, thereby slowing disease spread and limiting the overall pandemic size.
+Classical ODE-based compartmental models cannot capture this mechanism.
+We prove that for the same basic reproduction number $\mathcal{R}_{0}$, a mobility-based SIR model predicts a smaller final pandemic size compared to a classical SIR model, addressing the common overestimation issue in the literature.
+Another key contribution of this paper is the inference of mobility distribution from real-life data.
+For accurate forecasts and to match observed data, it is crucial to use a mobility distribution from real-world measurements.
+However, obtaining reliable mobility data is challenging due to large population sizes, privacy constraints, and inherent variability in individual movement patterns.
+We propose a method to infer the mobility distribution from accessible pandemic data, specifically the ratio of the infected population $I(t)$.
+We prove a one-to-one correspondence between the mobility distribution and the time series of $I(t)$, implying that the inference problem admits a unique solution.
+We also construct a training set of mobility distributions using Gaussian mixtures and employ an artificial-neural-network approach to identify the mobility distribution from real-life data.
 Our paper proceeds as follows.
-In section 2, we propose a mobility-based heterogeneous compartmental model, highlighting its distinctions from homogeneous models in terms of $\mathcal{R}_{0}$ and the final pandemic size.
-In section 3, we present an inverse problem aimed at inferring the mobility distribution from the time series of the infected population.
-We provide sufficient conditions to infer the mobility distribution uniquely and propose a machine-learning approach for inferring the mobility distribution.
+In section 2, we introduce a mobility-based compartmental model and emphasize its distinctions from homogeneous models, particularly in terms of the basic reproduction number and the final pandemic size.
+In section 3, we formulate the inverse problem to infer the mobility distribution from the time series of the infected population.
+We prove this inverse problem has a unique solution and propose a machine-learning approach to infer the mobility distribution.
 We conclude in section 4.
-2 Model description and properties
+2 Model description and epidemiological parameters
 2.1 A mobility-based SIRS model
-The baseline SIRS model delineates infectious disease dynamics within a population, categorizing individuals into susceptible (S), infected (I), and recovered (R) compartments.
-Consider a system with $N$ agents, in which agent $i$ holds a state $X_{i}$ that takes a value S, I, or R.
-We define $S(t)$, $I(t)$, and $R(t)$ as the ratios of agents in the three states as
-$S(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{S}\}},\quad I(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{I}\}},\quad R(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{R}\}}.$ (1)
-The baseline SIRS model posits that individuals transition cyclically between S, I, and R states through infection, recovery, and subsequent loss of immunity, all modeled as independent Poisson-point processes with constant rates.
-The well-known SIRS model describes the population dynamics of the three compartments as
-$\dot{S}(t)$ $=-{\beta}S(t)I(t)+\varepsilon R(t)$ (2)
-$\dot{I}(t)$ $={\beta}S(t)I(t)-\gamma I(t)$
+The baseline SIRS model categorizes agents into three compartments: susceptible (S), infected (I), and recovered (R).
+Consider a system with $N$ agents, in which agent $i$ holds a disease state $X_{i}$ that takes a value of S, I, or R.
+We define $S(t)$, $I(t)$, and $R(t)$ as the ratios of agents in the three states:
+$S(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{S}\}},\quad I(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{I}\}},\quad R(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{R}\}},$ (1)
+where $\mathbf{1}$ is the indicator function.
+The baseline SIRS model assumes individuals cycling between the S, I, and R states through infection, recovery, and loss of immunity.
+The baseline SIRS model without birth and death reads as
+$\dot{S}(t)$ $=-{\beta}S(t)I(t)+\varepsilon R(t),$ (2)
+$\dot{I}(t)$ $={\beta}S(t)I(t)-\gamma I(t),$
 $\dot{R}(t)$ $=\gamma I(t)-\varepsilon R(t),$
 where $\beta$ is the infection rate, $\gamma$ is the recovery rate, and $\varepsilon$ is the loss of immunity rate.
-However, this baseline model overlooks the heterogeneity among individuals within a population, which is critical for understanding disease transmission.
-Among various factors, mobility plays a pivotal role in epidemics by accelerating infection movement across regions through travel, commuting, and migration.
-We propose a mobility-based SIRS model where each agent $i$ has a time-independent mobility $m_{i}\in(0,1)$ that determines the infection and transmission rates of agent $i$.
-All S-I pairs (e.g., a pair of agents $i$ and $j$ with $X_{i}=\text{S}$ and $X_{j}=\text{I}$) are subject to internal infections, with the probability of agent $j$ infecting agent $i$ given by $\beta m_{i}m_{j}$.
-Infected agents recover at a constant rate $\gamma$, and recovered agents lose immunity at a constant rate $\varepsilon$, consistent with the baseline model (2).
+Mobility is a key factor in epidemics, influencing the spread of infection across regions through travel, commuting, and migration.
+We let each agent have a time-independent mobility $m_{i}\in(0,1)$, which determines their infection and transmission rates.
+For any S-I pair (i.e., a susceptible agent $i$ and an infected agent $j$), the probability of infection is given by $\beta m_{i}m_{j}$.
+Infected agents recover at a constant rate $\gamma$, and recovered agents lose immunity at a constant rate $\varepsilon$.
 Figure 1 illustrates the transitions between the three compartments and their respective rates.
@@
 We describe the dynamics using three population density functions of mobility—$S(m,t)$, $I(m,t)$, $R(m,t)$—associated with the susceptible, infected, and recovered compartments, respectively.
-Here, $S(m,t)\mathrm{d}m,$ represents the ratio of susceptible agents with mobility in the interval $[m,m+\mathrm{d}m,)$.
+Here, $S(m,t)\mathrm{d}m$ represents the ratio of susceptible agents with mobility in the interval $[m,m+\mathrm{d}m)$.
+Let $f$ be the mobility distribution function of the whole population and we assume $f$ is time-independent.
 Due to population conservation, we have
-$S(m,t)+I(m,t)+R(m,t)=f(m),\quad\text{for all}\leavevmode\nobreak\ t\geq 0,$ (3)
-where $f(m)$ denotes the time-independent mobility distribution function. The corresponding mobility-based SIRS model is
-$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},+\varepsilon R(m,t)$ (4a)
-$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},-\gamma I(m,t)$ (4b)
+$S(m,t)+I(m,t)+R(m,t)=f(m),\quad\text{for all}\leavevmode\nobreak\ t\geq 0.$ (3)
+While individuals with different mobility can interact in numerous ways, we assume that interactions between susceptible and infected individuals still follow the mass-action principle, as in the classical SIR model.
+We assume that the probability of infection between an S-I pair is proportional to the product of their mobilities.
+Therefore, the corresponding mobility-based SIRS model is
+$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}m\bar{m}S(m,t)I(\bar{m},t)\leavevmode\nobreak\ \mathrm{d}\bar{m}+\varepsilon R(m,t),$ (4a)
+$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}m\bar{m}S(m,t)I(\bar{m},t)\leavevmode\nobreak\ \mathrm{d}\bar{m}-\gamma I(m,t),$ (4b)
 $\frac{\partial}{\partial t}R(m,t)$ $=\gamma I(m,t)-\varepsilon R(m,t).$ (4c)
-In the mobility-based SIRS model, a susceptible agent with mobility $m$ can be infected by any infected agent with mobility $m^{\prime}$, leading to an integral representation of the internal infection process.
+In the mobility-based SIRS model, a susceptible agent with mobility $m$ can be infected by any infected agent with mobility $\bar{m}$, leading to an integral representation of the internal infection process.
 In our model, each agent’s mobility is time-independent, resulting in local terms for the recovery and loss of immunity processes.
@@
 2.2 Dynamics of the mobility-based SIRS model
-In Figure 2, we illustrate the time evolution of the three population densities in the mobility-based SIRS model (4) during the first infection wave.
+In Figure 2, we illustrate the time evolution of the mobility-based SIRS model (4) during the first infection wave.
 Initially, the population is evenly distributed across the mobility domain $[0,1]$.
@@
 Over time, the infected population recovers, resulting in an increased density in $R(m,t)$.
-This infection-recovery process occurs earlier in the high-mobility region compared to the low-mobility region.
+This infection-recovery process occurs sooner in the high-mobility region compared to the low-mobility region.
 (a) $S(m,t)$
@@
 (c) $R(m,t)$
-Figure 2: Evolution dynamics of the mobility-based SIRS model (3), with parameter $\beta=0.1,\varepsilon=0.01$, $\gamma=0.3$ and initial $S(m,0)=0.99,I(m,0)=0.01,R(m,0)=0$, for (a) the susceptible population $S(m,t)$, (b) the infected population $I(m,t)$, and (c) the recovered population $R(m,t)$.
+Figure 2: Evolution dynamics of the mobility-based SIRS model (3), with parameter $\beta=0.1,\varepsilon=0.01$, $\gamma=0.3$ and initial $S(m,0)=0.99,I(m,0)=0.01,R(m,0)=0$ for (a) the susceptible population $S(m,t)$, (b) the infected population $I(m,t)$, and (c) the recovered population $R(m,t)$.
 The initial mobility distribution plays a crucial role in determining the mobility heterogeneity of the population as well as shaping the aggregated dynamics.
@@
 $\langle g\rangle(t):=\int_{0}^{1}g(m,t)\leavevmode\nobreak\ \mathrm{d}m,$ (5)
-and illustrate the time evolution of the ratios of three compartments in Figure 3.
-We fix the second-order moment of the initial mobility distribution (i.e., $\langle m^{2}f\rangle$) in all simulations, which, as we will show in section 2.3, is proportional to the basic reproduction number $\mathcal{R}_{0}$.
-In addition, we use the initial conditions
+and illustrate the time evolution of the ratios of three compartments in Figure 3. We use the initial conditions
 $S(m,0)=(1-I_{0})f(m),\quad I(m,0)=I_{0}f(m),\quad R(m,0)=0,$ (6)
 which we refer to as the proportional initial condition.
-We will use this initial condition by default without further explanation.
-We choose three initial mobility distributions that all have the same second-order moment $\langle m^{2}f\rangle$, where $f_{1}$ and $f_{2}$ are restricted normal distributions on $[0,1]$ with
+To demonstrate the effect of heterogeneous mobility, we choose three initial mobility distributions $f_{1}$, $f_{2}$, and $f_{3}$, where $f_{1}$ and $f_{2}$ are restricted normal distributions on $[0,1]$ with
 $f_{1}\propto 0.8\mathcal{N}(0.1,0.02)+0.2\mathcal{N}(0.95,0.02),\quad f_{2}\propto\mathcal{N}(0.2,0.365),$ (7)
-and $f_{3}$ is a Direct delta function $f(m)=\delta(m-0.435)$.
+and $f_{3}$ is a Direct delta function $f_{2}(m)=\delta(m-0.435)$.
+All three mobility distributions have the same second-order moment $\langle m^{2}f\rangle$, which, as we will show in section 2.3, is proportional to the basic reproduction number.
+The mobility distribution $f_{1}$ is a mixture of two restricted Gaussian distributions and has two bumps centered at $m=0.1$ and $m=0.95$, indicating a polarized distribution of mobility among the population.
+The mobility distribution $f_{3}$ corresponds to the baseline SIRS model (2) in which everyone has the same mobility.
 (a) $\langle S\rangle(t)$
@@
 […]
+We observe from Figure 3 that at the end of the pandemic, over $70\%$ of the population are infected for the homogeneous population case $f_{3}$.
+In contrast, $f_{1}$ has polarized mobility distribution with a relatively small group of high-mobile individuals.
+When high-mobile individuals recover from infection, they can provide protection to uninfected low-mobile groups.
+Figure 3 shows that the final pandemic size regarding $f_{1}$ is only about $1/3$ of that of the baseline model.
+We derive and discuss the final pandemic size in detail in section 2.5.
 2.3 Basic reproduction number
 The basic reproduction number $\mathcal{R}_{0}$ is a key epidemiological metric that represents the average number of secondary infections produced by a single infected individual in a fully susceptible population.
-In this section, we derive $\mathcal{R}_{0}$ of the mobility-based SIRS model (4).
-We consider the next-generation matrix method [4] and discretize $I(m,t)$ and $S(m,t)$ on the grids $\{m_{i}\}_{i=1}^{M}$.
-We denote the discretized quantities by two vectors $\bm{x},\bm{y}\in\mathbb{R}^{M}$, in which
-$x_{i}(t)=I(m_{i},t),\quad y_{i}(t)=S(m_{i},t).$ (8)
-We consider a numerical quadrature to replace the integral in (4) with
-$\int_{0}^{1}m^{\prime}I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},\approx\sum_{i=1}^{M}w_{i}m_{i}x_{i},$ (9)
-where $w_{i}$ is the weight of the numerical quadrature at the grid point $m_{i}$. We write (4) in a discrete format and obtain
-$\dot{\bm{x}}=\bm{F}(\bm{x},\bm{y})-\bm{V}(\bm{x},\bm{y}),$ (10)
-where $\bm{F}$ is the rate of new infections from the susceptible population and $\bm{V}$ is the rate of loss of infections due to recovery. They are
-$F_{i}=\beta m_{i}y_{i}\sum_{j}w_{j}m_{j}x_{j},\quad V_{i}=\gamma x_{i}.$ (11)
-Consider the disease-free equilibrium, in which the whole population only contains susceptible individuals and its population density satisfies $S(m,t)\equiv f(m)$.
-Therefore, the disease-free equilibrium vectors $\bm{x}^{0},\bm{y}^{0}\in\mathbb{R}^{M}$ satisfy $x^{0}_{i}=f(m_{i})$ and $y^{0}_{i}=0$.
-We compute the Jacobi matrices of $\bm{F}$ and $\bm{V}$ with respect to $\bm{x}$ at the disease-free equilibrium vectors $\bm{x}^{0},\bm{y}^{0}$ and obtain
-$\frac{\partial\bm{F}}{\partial\bm{x}}(\bm{x}^{0},\bm{y}^{0})_{ij}=\beta m_{i}y^{0}_{i}m_{j}w_{j},\quad\frac{\partial\bm{V}}{\partial\bm{x}}(\bm{x}^{0},\bm{y}^{0})_{ij}=\gamma\delta_{ij},$ (12)
-where $\delta_{ij}$ is the Kronecker delta function. We further compute the next-generation matrix
-$\frac{\partial\bm{F}}{\partial\bm{x}}\left(\frac{\partial\bm{V}}{\partial\bm{x}}\right)^{-1}=\frac{\beta}{\gamma}\begin{pmatrix}m_{1}y^{0}_{1}\\ \vdots\\ m_{M}y^{0}_{M}\end{pmatrix}\begin{pmatrix}m_{1}w_{1}&\cdots&m_{M}w_{M}\end{pmatrix}_{j}.$ (13)
-We compute the largest eigenvalue of the next-generation matrix and obtain the basic reproduction number $\mathcal{R}_{0}=\beta\sum_{i}w_{i}m_{i}^{2}y^{0}_{i}/\gamma$.
-Notice that $w_{i}$ are the weights of grid points in the numerical quadrature, so as the number of grids $M$ goes to infinite, $\mathcal{R}_{0}$ converges to the integral
-$\mathcal{R}_{0}=\frac{\beta}{\gamma}\int_{0}^{1}m^{2}f(m)\leavevmode\nobreak\ \mathrm{d}m,.$ (14)
-The above derivations can be made rigorous by using the Fréchet derivative with an appropriate definition of the function space.
-When the entire population has a constant mobility $m_{0}$, equivalent to the baseline SIRS model (2), the basic reproduction number in (14) is
+In this section, we derive the basic reproduction number $\mathcal{R}_{0}$ of the mobility-based SIRS model (4) by using the next-generation operator (i.e., threshold operator) framework [25], which is essentially the infinite-dimensional version of the next generation matrix method [8].
+Recall the definition of $\mathcal{R}_{0}$ in [25]. For a generic infinite-dimensional SIRS model, suppose the infected population is defined on a compact set $\Omega$ and satisfies
+$\frac{\partial I(x,t)}{\partial t}=\int_{\Omega}h(x,y,S(x,t),I(y,t))\leavevmode\nobreak\ \mathrm{d}y-v(x)I(x,t),$ (8)
+where $h(x,y,S,I)$ is an incidence function and $v(x)$ is the recovery rate function. We define the threshold operator $\mathcal{L}$ as
+$\mathcal{L}u(x)=\int_{\Omega}\frac{u(y)}{v(x)}\left.\frac{\partial h(x,y,S,I)}{\partial I}\right|_{S=N(x),I=0}\leavevmode\nobreak\ \mathrm{d}y,$ (9)
+where $N(x)$ is the total population density, then the basic reproduction number $\mathcal{R}_{0}$ is the spectral radius of $\mathcal{L}$, which means
+$\mathcal{R}_{0}=\lim_{n\rightarrow\infty}\|(\mathcal{L})^{n}\|^{1/n}.$ (10)
+We refer [25] for detailed assumptions on $h$ and $v$.
+For the mobility-based SIRS model defined in (4), the total population density is $f(m)$ and the threshold operator $\mathcal{L}^{f}$ is
+$\mathcal{L}^{f}u(m)=\frac{\beta}{\gamma}mf(m)\int_{0}^{1}\bar{m}u(\bar{m})\mathrm{d}\bar{m}.$ (11)
+Therefore, the basic reproduction number $\mathcal{R}_{0}=\lim_{n\rightarrow\infty}\|(\mathcal{L}^{f})^{n}\|^{1/n}$, and we compute $\mathcal{R}_{0}$ in Theorem 1.
+Theorem 1.
+The basic reproduction number $\mathcal{R}_{0}$ of the mobility-based SIRS model (4) is
+$\mathcal{R}_{0}=\frac{\beta}{\gamma}\langle m^{2}f\rangle=\frac{\beta}{\gamma}\int_{0}^{1}m^{2}f(m)\leavevmode\nobreak\ \mathrm{d}m.$ (12)
+Proof.
+Recall the definition of the threshold operator $\mathcal{L}^{f}$ in (11).
+For any function $u(m)$, the image $\mathcal{L}^{f}u(m)$ lies in a 1-dimensional function space spanned by $mf(m)$.
+Consequently, $\mathcal{L}^{f}$ is a finite-rank operator, and thus, it must be compact.
+Therefore, the spectral radius of $\mathcal{L}^{f}$ is its maximal eigenvalue.
+Because $\mathcal{L}^{f}$ is a rank-one operator, $\mathcal{L}^{f}$ only has 1 nonzero eigenvalue and this nonzero eigenvalue corresponds to the eigenfunction $mf(m)$.
+We compute this eigenvalue by applying $\mathcal{L}^{f}$ to $mf(m)$ and obtaining
+$\mathcal{L}^{f}(mf(m))=\frac{\beta}{\gamma}mf(m)\int_{0}^{1}\bar{m}^{2}f(\bar{m})\mathrm{d}\bar{m}=\frac{\beta}{\gamma}\langle m^{2}f\rangle mf(m).$ (13)
+Therefore, the leading eigenvalue of $\mathcal{L}^{f}$ is $\frac{\beta}{\gamma}\langle m^{2}f\rangle$ and it is also the basic reproduction number $\mathcal{R}_{0}$.
+The basic reproduction number, $\mathcal{R}_{0}$, is a key dimensionless epidemiological parameter that represents the average number of secondary cases generated by one primary case.
+When $\mathcal{R}_{0}>1$, the disease-free equilibrium becomes unstable, allowing the disease to persist in the population.
+Conversely, if $\mathcal{R}_{0}<1$, the disease-free equilibrium is asymptotically stable, leading to the eventual extinction of the disease.
+The stability of both disease-free and endemic equilibria has been extensively studied over the past decades.
+For the original next-generation matrix method applied to finite-dimensional compartmental models, see [8, 7], and for various infinite-dimensional generalizations, refer to [24, 25, 22].
+By applying Theorems 3.16 and 3.17 in [24], one can readily show that the disease-free equilibrium $(f(m),0,0)$ of our model (4) is unstable when $\mathcal{R}_{0}>1$ and asymptotically stable when $\mathcal{R}_{0}<1$.
+When the entire population has a constant mobility $m_{0}$, equivalent to the baseline SIRS model (2), the basic reproduction number in (12) becomes
 $\mathcal{R}_{0}=\frac{\beta m_{0}^{2}}{\gamma}=\frac{\beta_{\text{eff}}}{\gamma},$ (14)
-which is consistent with the results of the baseline SIRS model (2).
-Let $\text{var}(f)$ denote the variance of a mobility distribution $f$. Assuming that all distributions that have the same $\mathcal{R}_{0}$, from (14) we have
+which is the same as the basic reproduction number of the baseline SIRS model (2) with the infection rate $\beta_{\text{eff}}$.
+Let $\text{var}(f)$ denote the variance of a mobility distribution $f$.
+In practice, we often estimate $\mathcal{R}_{0}$ by considering the early growth rate of cases and performing contact tracing studies, which is independent of models.
+Assuming that all distributions that have the same $\mathcal{R}_{0}$, from (12) we have
 $\langle mf\rangle^{2}=\frac{\gamma}{\beta}\mathcal{R}_{0}-\text{var}(f)\leq\frac{\gamma}{\beta}\mathcal{R}_{0}=\langle mf_{\delta}\rangle^{2}.$ (15)
-where $f_{\delta}$ is the Dirac delta distribution that has the same second-order moment as other distributions.
-Therefore, if we fix $\mathcal{R}_{0}$, the Dirac delta distribution $f_{\delta}$ maximizes the mean mobility $\langle mf\rangle$.
-In practice, we often estimate $\mathcal{R}_{0}$ by considering the early growth rate of cases and performing contact tracing studies, which is independent of model selection.
-As a result, using the value of $\mathcal{R}_{0}$ to estimate the pandemic size with the baseline SIRS model tends to result in overestimation.
+where $f_{\delta}$ is the Dirac delta distribution that has the same second-order moment as $f$.
+Therefore, the mean mobility $\langle mf_{\delta}\rangle$ of the Dirac delta distribution $f_{\delta}$ is larger than that of any other distribution $f$.
+Once $\mathcal{R}_{0}$ is determined first (using a model-free method), the homogeneous model (e.g., the baseline SIRS model) overestimates the mean mobility, which consequently leads to an overestimation of the final pandemic size.
 2.4 Mean mobility
-Mobility plays a crucial role in the mobility-based compartmental models.
-We examine the mean mobility of three populations.
-We focus on the mobility-based SIR model, which is defined by the following equations
-$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},$ (17a)
-$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},-\gamma I(m,t)$ (17b)
-$\frac{\partial}{\partial t}R(m,t)$ $=\gamma I(m,t).$ (17c)
-We define $Q_{k}^{\varphi}$ as the $k$th moment with respect to a nonnegative function $\varphi(m,t)$, which is
+Mobility plays a crucial role in the mobility-based compartmental models. We focus on the mobility-based SIR model, which is defined by the following equations,
+$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}m\bar{m}S(m,t)I(\bar{m},t)\leavevmode\nobreak\ \mathrm{d}\bar{m},$ (16a)
+$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}m\bar{m}S(m,t)I(\bar{m},t)\leavevmode\nobreak\ \mathrm{d}\bar{m}-\gamma I(m,t),$ (16b)
+$\frac{\partial}{\partial t}R(m,t)$ $=\gamma I(m,t),$ (16c)
+and examine the mean mobility of the susceptible compartment. Let $Q_{k}^{\varphi}$ be the $k$th moment with respect to a nonnegative function $\varphi(m,t)$, which means
 $Q_{k}^{\varphi}(t)=\int_{0}^{1}m^{k}\varphi(m,t)\leavevmode\nobreak\ \mathrm{d}m.$ (17)
+Under this definition, the ratio of the susceptible population is $Q_{0}^{S}(t)$, and the mean mobility of the susceptible compartment is $Q_{1}^{S}(t)/Q_{0}^{S}(t)$.
+We show the monotone property associated with the susceptible population in Theorem 2.
 Theorem 2.
-Let ${Q_{k}^{S}}(t)$ be the $k$th moment of $S(m,t)$ which is the solution of the mobility-based SIR model (17).
-If $f$ is not a Dirac delta distribution, then for any $k\geq 0$, the ratio ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases with respect to $t$.
+Let ${Q_{k}^{S}}(t)$ be the $k$th moment of $S(m,t)$ in the mobility-based SIR model (16).
+If $f$ is not a Dirac delta distribution, then for all $k\geq 0$, the ratio ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases with respect to time $t$.
 Proof.
-We multiply $m^{k}$ to (17a) and integrate $m$ over $[0,1]$. We obtain
+By multiplying $m^{k}$ to (16a) and integrating $m$ over $[0,1]$, we obtain that
 $\dot{Q}_{k}^{S}=-\beta Q_{1}^{I}Q_{k+1}^{S},\quad\text{for all}\leavevmode\nobreak\ k\geq 0.$ (18)
-By using (19), we obtain the time derivative of ${Q_{k+1}^{S}}/{Q_{k}^{S}}$
-$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{Q_{k+1}^{S}}{Q_{k}^{S}}\right)=\frac{1}{Q_{k}^{S}Q_{k}^{S}}\left(Q_{k}^{S}\dot{Q}_{k+1}^{S}-\dot{Q}_{k}^{S}Q_{k+1}^{S}\right)=-\frac{\beta Q_{1}^{I}}{Q_{k}^{S}Q_{k}^{S}}\left(Q_{k}^{S}Q_{k+2}^{S}-Q_{k+1}^{S}Q_{k+1}^{S}\right).$ (20)
-From Hölder’s inequality, we know that
-$Q_{k}^{S}Q_{k+2}^{S}-Q_{k+1}^{S}Q_{k+1}^{S}>0,$ (21)
-which implies that the time derivative in (20) is negative, so that ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases. ∎
+Using (18), we obtain the time derivative of ${Q_{k+1}^{S}}/{Q_{k}^{S}}$
+$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{Q_{k+1}^{S}}{Q_{k}^{S}}\right)=\frac{1}{(Q_{k}^{S})^{2}}\left(Q_{k}^{S}\dot{Q}_{k+1}^{S}-\dot{Q}_{k}^{S}Q_{k+1}^{S}\right)=-\frac{\beta Q_{1}^{I}}{(Q_{k}^{S})^{2}}\left[Q_{k}^{S}Q_{k+2}^{S}-(Q_{k+1}^{S})^{2}\right].$ (19)
+From Hölder’s inequality, we have that
+$Q_{k}^{S}Q_{k+2}^{S}-(Q_{k+1}^{S})^{2}>0,$ (20)
+which implies that the time derivative in (19) is negative, so that ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases over time $t$. ∎
 For $k=0$, the ratio in Theorem 2 simplifies to $Q_{1}^{S}(t)/Q_{0}^{S}(t)$, representing the mean mobility of the susceptible population.
@@
 However, this monotonic property does not necessarily apply to the infected and recovered populations, as illustrated in Figure 4.
-In addition, if $f$ is a Dirac delta distribution, then the mean mobility remains constant and is independent of time $t$.
+In addition, if $f$ is a Dirac delta distribution, then the entire population has a constant mobility, causing the mean mobility to remain unchanged over time.
 (a) $Q_{1}^{S}/Q_{0}^{S}$
@@
 (c) $Q_{1}^{R}/Q_{0}^{R}$
-[…]
+Figure 4: Time evolution of the mean mobility of three compartments in the mobility-based SIR model (16) with parameters $\beta=1.5$, $\gamma=0.13$, $I_{0}=0.0001$, and $f=f_{1}$ in (7).
+The mean mobility of the susceptible population decreases strictly over time, while the mean mobility of the infected and the recovered populations increases first and then decreases as time progresses.
 2.5 Final pandemic size
-The final pandemic size refers to the total number of individuals infected throughout an epidemic and is a crucial quantity to assess disease transmission dynamics.
-For the classical SIR model with a basic reproduction number $\mathcal{R}_{0}$, it is well known that the final pandemic size $R_{\infty}$ satisfies the equation
+The final pandemic size, defined as the total number of individuals infected over the course of an epidemic, is a critical quantity for assessing disease transmissions.
+For the classical SIR model, it is well known that the final pandemic size $R_{\infty}$ satisfies
 $R_{\infty}+\exp(-\mathcal{R}_{0}R_{\infty})=1,$ (21)
-if the initial infected population $I\approx 0$.
-In the real world, the population exhibits heterogeneous mobility, with high-mobility individuals likely being infected first during the early pandemic stages.
-Once recovered, they provide some protection to low-mobility individuals.
-In practice, $\mathcal{R}_{0}$ is typically estimated from the initial exponential growth rate, which is biased toward the infection rate of high-mobility individuals, thereby resulting in an overestimation of the final pandemic size.
+where $\mathcal{R}_{0}$ is the basic reproduction number.
+Equation (21) often overestimates the final pandemic size.
+For instance, when $\mathcal{R}_{0}=2$, Equation (21) suggests that $R_{\infty}$ is around $0.8$, implying that nearly $80\%$ of the population will ultimately be infected, which is a scenario rarely observed in practice.
+In reality, populations have heterogeneous mobility.
+High-mobility individuals are more likely to be infected early in a pandemic, and once they recover, they provide indirect protection to low-mobility individuals.
+Typically, $\mathcal{R}_{0}$ is estimated from the initial exponential growth rate, which is skewed by the higher infection rates among high-mobility individuals.
+This bias leads to an overestimation of the final pandemic size.
+Take the recent COVID-19 pandemic as an example.
+Studies estimate the basic reproduction number $\mathcal{R}_{0}$ for COVID-19 to be between 2 and 5 [16, 29, 6] in early 2020.
+However, later variants, such as Delta and Omicron, prove significantly more contagious.
+Studies suggest that the $\mathcal{R}_{0}$ of the Omicron variant is as high as 10 [27, 19, 2], which implies, according to the final pandemic size estimation in (21), that nearly the entire population would eventually become infected.
+However, empirical data indicate otherwise.
+As of December 2021, after two years of multiple infection waves caused by various SARS-CoV-2 variants, the seroprevalence in the United States is only 33.5%.
+Moreover, the largest infection wave, caused by the Omicron variant from December 2021 to February 2022, infects less than one-quarter of the total population [4].
+These numbers are far below the final pandemic size suggested by the classical SIR model.
 In this section, we derive the final pandemic size of the mobility-based SIR model (16) with the proportional initial condition (6).
-We rigorously demonstrate why the classic SIR model tends to overestimate the final size of pandemics, assuming the same $\mathcal{R}_{0}$ as in the mobility-based SIR model.
+We provide a quantitative explanation of why the classical SIR model tends to overestimate the final pandemic size in comparison to our mobility-based SIR model (16).
 Theorem 3.
-Let $R_{\infty}$ be the final pandemic size of the mobility-based SIR model (17) with the proportial initial condition (6). Assume $I=0\approx 0$, then $R_{\infty}$ satisfies
-$R_{\infty}+\langle f\exp\left(-\alpha m\beta/\gamma\right)\rangle$ $=1,$ (23)
-where $\alpha$ is the solution of $\alpha+\langle mf\exp\left(-\alpha m\beta/\gamma\right)\rangle=\langle mf\rangle$ on the interval $[0,1]$.
+Let $R_{\infty}$ be the final pandemic size of the mobility-based SIR model (16) with the proportial initial condition (6). Then the final pandemic size $R_{\infty}$ satisfies
+$R_{\infty}+\langle f\exp\left(-\frac{\alpha m\beta}{\gamma}\right)\rangle$ $=1,$ (22)
+where $\alpha$ is the root on the interval $[0,1]$ of the equation
+$\alpha+\langle mf\exp\left(-\frac{\alpha m\beta}{\gamma}\right)\rangle=\langle mf\rangle.$ (23)
 Proof.
-By solving (2) for $S(m,t)$, we obtain that
+We solve (16) for $S(m,t)$ and obtain that
 $S(m,t)=S(m,0)\exp\left(-\frac{\beta}{\gamma}m\langle mR\rangle(t)\right).$ (24)
@@
 $\langle mf\rangle-\langle mR\rangle(\infty)=\langle mS(m,0)\exp\left(-\frac{\beta}{\gamma}m\langle mR\rangle(\infty)\right)\rangle.$ (26)
-We use the initial condition (6) and assume that $S(m,0)\approx f(m)$, then $\alpha=\langle mR\rangle(\infty)$ satisfies
-$\alpha+\langle mf\exp\left(-\alpha m\beta/\gamma\right)\rangle=\langle mf\rangle.$ (27)
-Integrating (24) and taking $t$ to infinite, we obtain that $R_{\infty}=\langle R\rangle(\infty)$ satisfies (23). ∎
-Theorem 2 offers a method to compute the final pandemic size without the need to simulate the model (17) over time.
-The process involves first determining $\alpha$ by finding the root of (27) within the interval $[0,1]$.
-Once $\alpha$ is obtained, we then solve (23) to find $R_{\infty}$.
-Additionally, once the parameters $\beta$ and $\gamma$ are fixed, the mobility distribution $f(m)$ solely determines both the final pandemic size $R_{\infty}$ and the basic reproduction number $\mathcal{R}_{0}$ as given in (14).
-We will demonstrate in Theorem 3 that among all distributions that yield the same basic reproduction number, the Dirac delta distribution $f_{\delta}$, which corresponds to a homogeneous mobility of the population, maximizes the final pandemic size.
+Using the initial condition (6) and $S(m,0)\approx f(m)$, we find that $\alpha=\langle mR\rangle(\infty)$ satisfies
+$\alpha+\langle mf\exp\left(-\frac{\alpha m\beta}{\gamma}\right)\rangle=\langle mf\rangle.$ (27)
+We integrate (24) over $m$ and take $t$ to infinite, then we obtain that $R_{\infty}=\langle R\rangle(\infty)$ satisfies (22). ∎
+Theorem 3 offers a method to compute the final pandemic size of the model (16) efficiently.
+We first determine $\alpha$ by finding the root of (27) within the interval $[0,1]$ and then use the value of $\alpha$ to compute $R_{\infty}$ using (22).
+Once the parameters $\beta$ and $\gamma$ are fixed, the mobility distribution $f(m)$ solely determines both the final pandemic size $R_{\infty}$ and the basic reproduction number $\mathcal{R}_{0}$ as given in (12).
+In Theorem 4, we prove that among all distributions that yield the same basic reproduction number $\mathcal{R}_{0}$, the Dirac delta distribution $f_{\delta}$, which corresponds to a population with the same mobility, maximizes the final pandemic size.
 Theorem 4.
-Let $R_{\infty}$ and $R^{\delta}_{\infty}$ be the final pandemic sizes associated with a distribution $f$ and a Dirac delta distribution $f_{\delta}$, respectively.
-Suppose that $f$ and $f_{\delta}$ yield the same basic reproduction number $\mathcal{R}_{0}$ (14), it holds that $R_{\infty}\leq R^{\delta}_{\infty}$ with the equality occurring only if $f=f_{\delta}$.
-Theorem 2 provides an approach to compute the final pandemic size. We introduce another perspective to represent the final pandemic size to prove Theorem 3.
+Let $R_{\infty}$ and $R^{\delta}_{\infty}$ be the final pandemic sizes associated with the mobility distribution $f$ and the Dirac delta mobility distribution $f_{\delta}$, respectively.
+Suppose that $f$ and $f_{\delta}$ yield the same basic reproduction number $\mathcal{R}_{0}$ (12), then $R_{\infty}\leq R^{\delta}_{\infty}$ and the equality occurs only if $f=f_{\delta}$.
+Theorem 4 elucidates why the classical SIR model tends to overestimate the final pandemic size compared to our mobility-based model.
+The classical SIR model assumes a homogeneous population, which corresponds to a Dirac delta mobility distribution $f_{\delta}$ in the framework of our model (16).
+Its final pandemic size $R_{\infty}^{\delta}$ is greater than that of any other mobility distribution, given that all distributions share the same basic reproduction number $\mathcal{R}_{0}$.
+We prepare the proof of Theorem 4 with the following lemma.
+Lemma 5.
+Let $I(m,t)$ be the mobility distribution of the infected population in the mobility-based SIR model (16) with the proportional initial condition (6). Then for all $t\geq 0$, it holds that
+$\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{\langle m^{2}f\rangle}{\langle mf\rangle},$ (28)
+and the equality occurs only if $f$ is a Dirac delta distribution.
 Proof.
-From (17), we have
-$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}t}=-\beta Q_{1}^{S}Q_{1}^{I},\quad\frac{\mathrm{d}Q_{0}^{R}}{\mathrm{d}t}=\gamma Q_{0}^{I},\quad t\in[0,\infty).$ (28)
-Since $Q_{0}^{R}(t)$ monotonely increases with respect to $t$, so we have a one-to-one correspondence between $Q_{0}^{R}(t)$ and $t$.
-Therefore, we can view all functions of $t$ (including $Q_{0}^{S}$) as a function of $Q_{0}^{R}$.
-Equation (28) becomes
-$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}Q_{0}^{R}}=-\frac{\beta}{\gamma}\widetilde{H}(Q_{0}^{R})Q_{0}^{S},\quad Q_{0}^{R}\in[Q_{0}^{R}(0),Q_{0}^{R}(\infty)),$ (29)
-where
-$\widetilde{H}(Q_{0}^{R})=H(t)=\frac{Q_{1}^{S}(t)Q_{1}^{I}(t)}{Q_{0}^{S}(t)Q_{0}^{I}(t)}.$ (30)
-Recall that $Q_{0}^{R}(0)=0$. We integrate (29) over the interval $[0,Q_{0}^{R}(\infty)]$ and obtain
-$Q_{0}^{S}(\infty)=\exp\left(-\frac{\beta}{\gamma}\int_{0}^{Q_{0}^{R}(\infty)}\widetilde{H}(\alpha)\leavevmode\nobreak\ \mathrm{d}\alpha\right).$ (31)
-Using $\lim_{t\to\infty}Q_{0}^{I}(t)=0$ and change of variable, we obtain
-$Q_{0}^{R}(\infty)+\exp\left(-\frac{\beta}{\gamma}\int_{0}^{\infty}H(t)\leavevmode\nobreak\ \mathrm{d}Q_{0}^{R}(t)\right)=1.$ (32)
-As $R_{\infty}^{\delta}$ satisfies
-$R^{\delta}_{\infty}+\exp\left(-\mathcal{R}_{0}R^{\delta}_{\infty}\right)=1,$ (33)
-it is sufficient to prove that
-$H(t)\leq\frac{\gamma}{\beta}\mathcal{R}_{0}=\langle m^{2}f\rangle,\quad\text{for all\leavevmode\nobreak\ }t>0,$ (34)
-where the equal sign holds only if $f=f_{\delta}$.
-It is not difficult to verify that the equality holds when $f=f_{\delta}$.
-We will assume that $f\neq f_{\delta}$ and prove that $H(t)<\langle m^{2}f\rangle$ in the following proof.
-From Theorem 1 for $k=1$ and Lemma 4, we have
-$\frac{Q_{1}^{S}(t)}{Q_{0}^{S}(t)}<\langle mf\rangle\quad\text{and}\quad\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{\langle m^{2}f\rangle}{\langle mf\rangle},\quad\text{for all\leavevmode\nobreak\ }t>0.$ (35)
-Combining it with (30), we complete the proof. ∎
-Lemma 4.
-Let $I(m,t)$ be the solution of the mobility-based SIR model (17) with the proportional initial condition (6). For all $t>0$, it holds that
-$\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{\langle m^{2}f\rangle}{\langle mf\rangle},$ (36)
-with equality occurring only if $f$ is a Dirac delta distribution.
-Proof.
-If $f$ is a Dirac delta distribution $f(m)=\delta(m-m_{0})$, then both sides of (36) equal $m_{0}$. We assume that $f$ is not a Dirac delta distribution in the following proof.
+If $f$ is a Dirac delta distribution $f(m)=\delta(m-m_{0})$, then both sides of (28) equal $m_{0}$. We assume that $f$ is not a Dirac delta distribution in the rest of the proof.
 Using (16), we compute the time derivative of ${Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}$ and obtain
 $\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{Q_{1}^{I}}{Q_{0}^{I}}\right)=\frac{\beta Q_{1}^{I}Q_{1}^{S}}{Q_{0}^{I}}\left(\frac{Q_{2}^{S}}{Q_{1}^{S}}-\frac{Q_{1}^{I}}{Q_{0}^{I}}\right).$ (29)
-We denote the instants that $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}\right)$ changes its sign as $t_{k}$ and partition the time interval $(0,\infty)$ into open intervals $\Omega_{k}$ with $t_{k}$ as indicated in Figure 5.
+We denote the time that $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}\right)$ changes its sign as $t_{k}$ and partition the time interval $(0,\infty)$ into open intervals $\Omega_{k}$ as shown in Figure 5.
 Figure 5: A partition of intervals due to the sign of (29).
@@
 $\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{Q_{1}^{I}}{Q_{0}^{I}}\right)(0)=\frac{\beta Q_{1}^{I}(0)Q_{1}^{S}(0)}{Q_{0}^{I}(0)}\left(\frac{\langle m^{2}f\rangle}{\langle mf\rangle}-\langle mf\rangle\right)>0,$ (30)
-which implies that $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}}/{Q_{0}^{I}}\right)$ is nonnegative on the interval $\Omega_{1}$.
-Due to the smoothness, the signs of (37) interlace on $\Omega_{1}$, $\Omega_{2}$, $\Omega_{3}$, $\cdots$.
-If $t\in\cup_{k=1}^{\infty}\overline{\Omega_{2k-1}}$, then we know (37) is nonnegative, which implies that
+so $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}}/{Q_{0}^{I}}\right)$ is positive on the interval $\Omega_{1}$.
+Due to the smoothness, the signs of (29) interlace on $\Omega_{1}$, $\Omega_{2}$, $\Omega_{3}$, $\ldots$.
+If $t\in\cup_{k=1}^{\infty}{\Omega}_{2k-1}$ or $t\in\{t_{1},t_{2},\ldots\}$, then we know (29) is nonnegative, which implies that
 $\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{Q_{2}^{S}(t)}{Q_{1}^{S}(t)}<\frac{Q_{2}^{S}(0)}{Q_{1}^{S}(0)}=\frac{\langle m^{2}f\rangle}{\langle mf\rangle},$ (31)
 where the second inequality uses Theorem 2 for the case $k=1$.
-If $t\in\cup_{k=1}^{\infty}\Omega_{2k}$ (supposing $t\in\Omega_{2k}$), then ${Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}$ decreases with respect to $t$ on the interval $\Omega_{2k}$.
+If $t\in\cup_{k=1}^{\infty}\Omega_{2k}$ (supposing $t\in\Omega_{2k}$), then ${Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}$ decreases strictly with respect to $t$ on the interval $\Omega_{2k}$.
 Due to continuity, we have
 $\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{Q_{1}^{I}(t_{2k-1})}{Q_{0}^{I}(t_{2k-1})}<\frac{\langle m^{2}f\rangle}{\langle mf\rangle},\quad\text{for all\leavevmode\nobreak\ }t\in\Omega_{2k-1},$ (32)
-where the second inequality holds from (39).
+where the second inequality holds because of (31). Equations (31) and (32) complete the proof. ∎
+Now, we turn to the proof of Theorem 4.
+Proof.
+From (16), we have
+$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}t}=-\beta Q_{1}^{S}Q_{1}^{I},\quad\frac{\mathrm{d}Q_{0}^{R}}{\mathrm{d}t}=\gamma Q_{0}^{I},\quad\text{for all\leavevmode\nobreak\ }t\geq 0.$ (33)
+Since $Q_{0}^{R}(t)$ increases strictly with respect to $t$, we have a one-to-one correspondence between $Q_{0}^{R}$ and $t$. Using the chain rule, we have
+$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}Q_{0}^{R}}\frac{\mathrm{d}Q_{0}^{R}}{\mathrm{d}t}=-\beta Q_{1}^{S}Q_{1}^{I},$ (34)
+yielding
+$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}Q_{0}^{R}}=-\frac{\beta}{\gamma}\frac{Q_{1}^{S}Q_{1}^{I}}{Q_{0}^{I}}.$ (35)
+Due to the one-to-one correspondence between $t$ and $Q_{0}^{R}$, we can view all functions of $t$ (including $Q_{0}^{S}$) as a function of $Q_{0}^{R}$ and rewrite (35) as
+$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}Q_{0}^{R}}=-\frac{\beta}{\gamma}\widetilde{H}(Q_{0}^{R})Q_{0}^{S},$ (36)
+where $\widetilde{H}$ is a function of $Q_{0}^{R}$ and defined by
+$\widetilde{H}(Q_{0}^{R}(t))=\frac{Q_{1}^{S}(t)Q_{1}^{I}(t)}{Q_{0}^{S}(t)Q_{0}^{I}(t)}.$ (37)
+Recall that $Q_{0}^{R}(0)=0$ and $Q_{0}^{R}(\infty)=R_{\infty}$. We integrate (36) over the interval $[0,R_{\infty}]$ and use $Q_{0}^{S}(\infty)=1-R_{\infty}$, obtaining
+$R_{\infty}+\exp\left(-\frac{\beta}{\gamma}\int_{0}^{R_{\infty}}\widetilde{H}(\alpha)\leavevmode\nobreak\ \mathrm{d}\alpha\right)=1.$ (38)
+Recall that $R_{\infty}^{\delta}$ satisfies (21). We rewrite (21) as
+$R^{\delta}_{\infty}+\exp\left(-\int_{0}^{R_{\infty}^{\delta}}\mathcal{R}_{0}\leavevmode\nobreak\ \mathrm{d}\alpha\right)=1.$ (39)
+From Theorem 2 for $k=1$ and Lemma 5, we have
+$\frac{Q_{1}^{S}(t)}{Q_{0}^{S}(t)}<\langle mf\rangle\quad\text{and}\quad\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}<\frac{\langle m^{2}f\rangle}{\langle mf\rangle},\quad\text{for all\leavevmode\nobreak\ }t>0.$ (40)
+Using (12), (37), and (40), we obtain that
+$\frac{\beta}{\gamma}\widetilde{H}(\alpha)\leq\mathcal{R}_{0},\quad\text{for all\leavevmode\nobreak\ }\alpha\in[0,R_{\infty}),$ (41)
+and the equality holds only if $f$ is a Dirac delta distribution (i.e., $f=f_{\delta}$).
+From (38),(39), and (41), we obtain that $R_{\infty}\leq R_{\infty}^{\delta}$ and the equality holds only if $f=f_{\delta}$. ∎
 3 Learning the mobility distribution from data
-Mobility holds critical significance in epidemic models as it directly influences the spread and dynamics of infectious diseases within populations.
-Understanding how people move between locations allows for more accurate predictions of disease transmission and enhances the precision of epidemic models to reflect real-world scenarios.
-However, the mobility of a population involves complex factors, including travel patterns, migration, and commuting, which are often impossible to measure directly in real-world situations.
-Realistically, the available data in epidemics is the number of infected individuals, which translates to the ratio of infected agents at time $t$:
+The mobility distribution plays a significant role in the mobility-based compartmental models.
+A realistic mobility distribution allows for more accurate predictions of disease transmission and enhances the precision of epidemiological models to reflect real-world scenarios.
+However, it is often impossible to measure the mobility of a whole population directly in real-world situations.
+In practice, the available data during epidemics is the number of infected individuals, which corresponds to the infected population ratio, defined as
 $Q_{0}^{I}(t)=\int_{0}^{1}I(m,t)\leavevmode\nobreak\ \mathrm{d}m,$ (42)
 in the mobility-based SIRS model (4).
-In this section, we infer the mobility distribution $f(m)\in\mathcal{L}^{1}(0,1)$ from the time series of $Q_{0}^{I}(t)$, offering both a theoretical guarantee of unique inference and a machine-learning-based numerical approach.
-3.1 Theoretical guarantees
+In this section, we propose an inverse problem to infer the mobility distribution $f(m)\in\mathcal{L}^{1}(0,1)$ from the time-dependent infected ratio $Q_{0}^{I}(t)$.
+We prove that $f(m)$ is identifiable from $Q_{0}^{I}(t)$ (i.e., the inverse problem has a unique solution) and provide a machine-learning-based approach to infer $f(m)$ from real data.
+3.1 Identifiability of the mobility distribution
+In this section, we show that the observation $Q_{0}^{I}(t)$ contains sufficient information to infer the mobility distribution $f(m)$. In other words, this inference problem has a unique solution.
 Theorem 6.
-Let $I(m,t)$ be the infected mobility distribution of the mobility-based SIRS model (4) with the initial condition (6), yielding the 0th-order moment $Q_{0}^{I}(t)$.
-Assume that $Q_{0}^{I}(m,t)$ is arbitrarily smooth (i.e., $Q_{0}^{I}\in\mathcal{C}^{\infty}(0,\infty)$) and all the parameters ($I_{0}$, $\beta$, $\gamma$, and $\varepsilon$) are fixed and not equal to 0, then there is a one-to-one correspondence between $f$ and $Q_{0}^{I}$.
-This theorem indicates that for any two different mobility distributions ($f\neq\tilde{f}$), the corresponding time series of the infected numbers ($Q_{0}^{I}$ and $Q_{0}^{\tilde{I}}$) also differ.
-This implies that it is sufficient to infer $f$ from $I$ theoretically and this inference problem has a unique solution.
-We prove this theorem at the end of this subsection.
-Recall the moment notation in (18). We multiply $m^{k}$ to (4) and obtain the equations for moments
-$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{S}$ $=-\beta Q_{1}^{I}Q_{k+1}^{S}+\varepsilon Q_{k}^{R}$ (42a)
-$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{I}$ $=\beta Q_{1}^{I}Q_{k+1}^{S}-\gamma Q_{k}^{I}$ (42b)
+Let $I(m,t)$ be the infected mobility distribution of the mobility-based SIRS model (4) with the initial condition (6) and $Q_{0}^{I}(t)$ be the 0th order moment of $I(m,t)$, defined in (42).
+Assume that $Q_{0}^{I}(t)$ is arbitrarily smooth (i.e., $Q_{0}^{I}(t)\in\mathcal{C}^{\infty}(0,\infty)$) and that all the parameters ($I_{0}$, $\beta$, $\gamma$, and $\varepsilon$) are fixed and not equal to 0, then there is a one-to-one correspondence between $f$ and $Q_{0}^{I}$.
+Recall the moment notation in (17). We multiply $m^{k}$ to (4) and integrate over $m$, obtaining the equations for moments
+$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{S}$ $=-\beta Q_{1}^{I}Q_{k+1}^{S}+\varepsilon Q_{k}^{R},$ (43a)
+$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{I}$ $=\beta Q_{1}^{I}Q_{k+1}^{S}-\gamma Q_{k}^{I},$ (43b)
 $\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{R}$ $=\gamma Q_{k}^{I}-\varepsilon Q_{k}^{R}.$ (43c)
-Consider another set of solutions $\tilde{S}(m,t),\tilde{I}(m,t),\tilde{R}(m,t)$ of (4) with the initial condition proportional to $\tilde{f}(m)$.
+Let $\tilde{S}(m,t),\tilde{I}(m,t),\tilde{R}(m,t)$ be the mobility distributions of (4) with the initial condition
+$\tilde{S}(m,0)=(1-I_{0})\tilde{f}(m),\quad\tilde{I}(m,0)=I_{0}\tilde{f}(m),\quad\tilde{R}(m,0)=0,$ (44)
+and $Q_{k}^{\tilde{S}},Q_{k}^{\tilde{I}},Q_{k}^{\tilde{R}}$ be their $k$-th order moments.
 Lemma 7.
 The following three statements are equivalent:
-$Q_{k}^{S}=Q_{k}^{\tilde{S}},\quad Q_{k}^{I}=Q_{k}^{\tilde{I}},\quad\text{and}\quad Q_{k}^{R}=Q_{k}^{\tilde{R}}.$ (43)
+$Q_{k}^{S}(t)=Q_{k}^{\tilde{S}}(t),\quad$ for all $\text{for all\leavevmode\nobreak\ }t\geq 0,$ (45)
+$Q_{k}^{I}(t)=Q_{k}^{\tilde{I}}(t),\quad$ for all $\text{for all\leavevmode\nobreak\ }t\geq 0,$
+$Q_{k}^{R}(t)=Q_{k}^{\tilde{R}}(t),\quad$ for all $\text{for all\leavevmode\nobreak\ }t\geq 0.$
 Proof.
-We add (42a) and (42b) and obtain
+We add (43a) and (43b) together and obtain
 $\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{S}+\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{I}=\varepsilon Q_{k}^{f}-(\varepsilon+\gamma)Q_{k}^{I}-\varepsilon Q_{k}^{S},$ (46)
-which utilizes the fact that $Q_{k}^{f}=Q_{k}^{S}+Q_{k}^{I}+Q_{k}^{R}$. Similarly, we have
+which uses the fact that $Q_{k}^{f}=Q_{k}^{S}+Q_{k}^{I}+Q_{k}^{R}$. Similarly, we have
 $\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{\tilde{S}}+\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{\tilde{I}}=\varepsilon Q_{k}^{\tilde{f}}-(\varepsilon+\gamma)Q_{k}^{\tilde{I}}-\varepsilon Q_{k}^{\tilde{S}}.$ (47)
 Step 1.
-We first show that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$ is equivalent to $Q_{k}^{S}=Q_{k}^{\tilde{S}}$.
+We show that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$ is equivalent to $Q_{k}^{S}=Q_{k}^{\tilde{S}}$.
 Suppose that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$.
@@
 $\frac{\mathrm{d}}{\mathrm{d}t}\left(Q_{k}^{S}-Q_{k}^{\tilde{S}}\right)=\varepsilon(Q_{k}^{f}-Q_{k}^{\tilde{f}})-\varepsilon(Q_{k}^{S}-Q_{k}^{\tilde{S}}).$ (48)
-Since $Q_{k}^{I}(0)=I_{0}Q_{k}^{f}$ and $Q_{k}^{\tilde{I}}(0)=I_{0}Q_{k}^{\tilde{f}}$, we use $Q_{k}^{I}(0)=Q_{k}^{\tilde{I}}(0)$ and obtain
-$Q_{k}^{f}=Q_{k}^{\tilde{f}},\quad Q_{k}^{S}(0)-Q_{k}^{\tilde{S}}(0)=S_{0}\left(Q_{k}^{f}-Q_{k}^{\tilde{f}}\right)=0.$ (47)
+Since $Q_{k}^{I}(0)=I_{0}Q_{k}^{f}$ and $Q_{k}^{\tilde{I}}(0)=I_{0}Q_{k}^{\tilde{f}}$, we use $Q_{k}^{I}(0)=Q_{k}^{\tilde{I}}(0)$ and obtain $Q_{k}^{f}=Q_{k}^{\tilde{f}}$, so that
+$Q_{k}^{S}(0)-Q_{k}^{\tilde{S}}(0)=(1-I_{0})\left(Q_{k}^{f}-Q_{k}^{\tilde{f}}\right)=0.$ (49)
 Therefore, $e(t)=Q_{k}^{S}(t)-Q_{k}^{\tilde{S}}(t)$ satisfies a linear ODE
@@
 We then show that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$ is equivalent to $Q_{k}^{R}=Q_{k}^{\tilde{R}}$.
-Suppose that $Q_{k}^{R}=Q_{k}^{\tilde{R}}$.
-Therefore, we have $\mathrm{d}Q_{k}^{R}/\mathrm{d}t=\mathrm{d}Q_{k}^{\tilde{R}}/\mathrm{d}t$.
+Suppose that $Q_{k}^{R}=Q_{k}^{\tilde{R}}$, then $\mathrm{d}Q_{k}^{R}/\mathrm{d}t=\mathrm{d}Q_{k}^{\tilde{R}}/\mathrm{d}t$.
 From (43c), we have $Q_{k}^{I}=Q_{k}^{\tilde{I}}$.
@@
 implying that $Q_{k}^{R}=Q_{k}^{\tilde{R}}$. ∎
-Theorem 5 and Lemma 6 indicate that there is a one-to-one correspondence between $f$ and $Q_{0}^{S}$ and that between $f$ and $Q_{0}^{R}$.
-This implies that, alternatively, we can also infer the mobility distribution $f$ from the time series of $Q_{0}^{S}$ or $Q_{0}^{R}$ if they are available and the inference problems yield unique solutions.
+Theorem 6 and Lemma 7 indicate that $f(m)$ also has a one-to-one correspondence to $Q_{0}^{S}(t)$ and $Q_{0}^{R}(t)$.
+This implies that, alternatively, we can infer the mobility distribution $f(m)$ from $Q_{0}^{S}(t)$ or $Q_{0}^{R}(t)$ if they are available, and the inference problems have unique solutions.
 Lemma 8.
-Let $L$ be a positive integer and the moments satisfy
+Let $L$ and $K$ be positive integers and $K\geq 2$. Assume the moments satisfy
 $\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{S,I,R}(0)$ $=\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad k=0,1,\ldots,K-1-\ell,$ (53)
-$\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{S,I,R}(0)$ $\neq\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad k=K-\ell.$
+$\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{S,I,R}(0)$ $\neq\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad k=K-\ell,$
 for $\ell=0,1,\ldots,L-1$, then (53) also holds for $\ell=L$.
@@
 Proof.
-We use the same moment notation (18). Since $f\neq\tilde{f}$, we can find a positive integer $K$ such that
+We use the same moment notation (17). Assume that $f\neq\tilde{f}$, then we can find a positive integer $K$ such that
 $Q_{k}^{f}=Q_{k}^{\tilde{f}},\quad\text{\leavevmode\nobreak\ for all}\leavevmode\nobreak\ k=0,1,\ldots,K-1,$ (58)
@@
 $Q_{K}^{S,I,R}(0)\neq Q_{K}^{\tilde{S},\tilde{I},\tilde{R}}(0).$
-We plug the initial values into (42) and obtain
+We plug the initial values of moments into (43) and obtain
 $\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{S,I,R}(0)=\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad\text{for all}\leavevmode\nobreak\ k=0,1,\ldots,K-2,$ (60)
@@
 By induction, we know (53) hold for $\ell=2,3,\ldots,K$.
-In particular, we use the inequality result in Lemma 7 for $\ell=K$ and $k=0$ yielding
+In particular, we use the inequality result in Lemma 8 for $\ell=K$ and $k=0$ and obtain
 $\frac{\mathrm{d}^{K}}{\mathrm{d}t^{K}}Q_{0}^{I}(0)\neq\frac{\mathrm{d}^{K}}{\mathrm{d}t^{K}}Q_{0}^{\tilde{I}}(0).$ (62)
 Therefore, we have $Q_{0}^{I}\neq Q_{0}^{\tilde{I}}$ and the correspondence between $f$ and $Q_{0}^{I}$ is one-to-one.
-3.2 Learning the mobility distribution using a neural-network framework
-Mobility distribution significantly impacts the pandemic size and transmission.
-However, in reality, it is not feasible to directly measure the mobility density functions of different compartments during pandemics.
-Realistically, the available data in epidemics is often limited to the number of infected individuals.
-Theorem 5 suggests a one-to-one correspondence between the mobility distribution $f(m)$ and the ratio of the infected population $\langle I\rangle(t)$.
-Thus, a natural approach is to infer the mobility distribution from the early-stage pandemic data to make accurate predictions for the future course of the pandemic.
-We propose a neural-network-based approach to learn the operator from the time-series $\langle I\rangle(\cdot)$ to the mobility distribution $f(\cdot)$.
-Assume that the pandemic data takes the form of $\mathbf{I}=(\langle I\rangle(t_{1}),\ldots,\langle I\rangle(t_{N_{t}}))^{T}$, which is a time series of the ratio of the infected population at discrete time steps.
-We discretize $\mathbf{f}$ using the discrete cosine transform (DCF) and denote the first $N_{f}$ modes as $\mathbf{f}=(\hat{f}_{1},\ldots,\hat{f}_{N_{f}})^{T}$.
-We consider a feedforward neural network $\mathrm{G}_{\theta}$ to learn the mapping between $\mathbf{I}$ and $\mathbf{f}$:
+3.2 Numerical inference of the mobility distribution
+Theorem 6 establishes a one-to-one correspondence between the mobility distribution $f(m)$ and the infected population ratio $Q_{0}^{I}(t)$.
+This provides a theoretical basis for inferring the mobility distribution from time series data of the infected population.
+One can infer the mobility distribution from early-stage pandemic data and use it to predict the future course of the pandemic.
+In this section, we propose a numerical approach using neural networks to achieve this inference.
+3.2.1 A neural-network-based learning framework
+Let $\langle I\rangle(t)$ be the ratio of the infected population at time $t$ (i.e., $\langle I\rangle(t)=\int_{0}^{1}I(m,t)\leavevmode\nobreak\ \mathrm{d}m$) and $\mathbf{I}$ be the observed time series of $\langle I\rangle(t)$, which takes the form of
+$\mathbf{I}=(\langle I\rangle(t_{1}),\langle I\rangle(t_{2}),\ldots,\langle I\rangle(t_{n_{t}}))^{T}.$ (63)
+Since the data is a finite-dimensional vector, we are only able to infer the mobility distribution $f$ in a finite-dimensional space.
+Let $\bar{f}$ be the spatial discretization of $f$ with $M$ uniform grids on the interval $[0,1]$.
+We take the discrete cosine transform (DCT) of $\bar{f}$ and denote the DCT of $\bar{f}$ by $\hat{f}=\sqrt{M}(1,\hat{f}_{1},\ldots,\hat{f}_{n_{f}},\ldots)$.
+Let $\mathbf{f}$ be the first $n_{f}$ nontrivial modes of $\hat{f}$, i.e.,
+$\mathbf{f}=(\hat{f}_{1},\hat{f}_{2},\ldots,\hat{f}_{n_{f}})^{T}.$ (64)
+We formulate the inference problem using a feedforward neural network $\mathrm{G}_{\theta}$ to learn the mapping from $\mathbf{I}$ to $\mathbf{f}$:
 $\mathbf{f}=\mathrm{G}_{\theta}(\mathbf{I}).$ (65)
-We sample the mobility distribution $f$ and obtain $\langle I\rangle$ by simulating the mobility-based SIRS model (4) to form the training set.
-When using neural networks to learn unknown parameters or functions, it is crucial to sample a training set that encompasses a wide range of scenarios.
-In our context, this involves properly sampling the mobility distribution comprehensively and robustly.
-Since a Gaussian mixture model is a universal approximator of probability densities, we sample $f$ by a Gaussian mixture with random coefficients.
+To build the training set, we sample various mobility distributions $f(m)$ and simulate the mobility-based SIRS model (4) to generate the corresponding time series of $\langle I\rangle(t)$.
+When using neural networks to learn unknown parameters or functions, it’s essential to sample a training set that covers a wide range of scenarios, ensuring the model is effective across the prediction regime.
+In our context, this involves sampling the mobility distribution comprehensively and robustly.
+Since the Gaussian mixture model is a universal approximator of probability densities [13], we sample $f$ as a Gaussian mixture model with random parameters.
 In particular, we let
-${f}\propto\sum_{n=1}^{3}r_{n}\mathcal{N}(\mu_{n},\sigma_{n}^{2}),$ (62)
-where $\mathcal{N}(\mu,\sigma^{2})$ is the Gaussian density function with mean $\mu$ and variance $\sigma^{2}$ and $r_{n}$, $\mu_{n}$, and $\sigma_{n}$ are all uniformly distributed on $[0,1]$, $[-0.2,1.2]$, and $[0.01,0.26]$, respectively.
-Suppose the dataset $\mathcal{D}$ contains $K$ pairs of $\mathbf{I}$ and $\mathbf{f}$, which we obtain by discretizing $\langle I\rangle$ and $f$.
-We parametrize $\mathrm{G}_{\theta}$ with a fully-connected feedforward neural network and minimize the loss function
-$\mathrm{L}(\theta)=\frac{1}{K}\sum_{i=1}^{K}\left[\mathrm{G}_{\theta}(\mathbf{I}_{i})-\mathbf{f}_{i}\right]^{2}.$
-In our simulations, we use a six-layer-fully connected neural network.
-The width of layers are 128, 512, 512, 512, 128, and 32, respectively.
-Each layer includes a Sigmoid activation function and $L^{2}$ regularization with a magnitude of $0.005$.
-In addition, we choose $N_{f}=8$ and $N_{t}=101$.
+${f}\propto\sum_{i=1}^{\alpha}w_{i}\mathcal{N}(\mu_{i},\sigma_{i}^{2})$ (66)
+where $\mathcal{N}(\mu,\sigma^{2})$ is the Gaussian probability density function and $\mu_{i}$, $\sigma_{i}$, and $w_{i}$ are random numbers.
+Suppose the dataset contains $K$ pairs of $\mathbf{I}_{k}$ and $\mathbf{f}_{k}$, which we discretize from $\langle I_{k}\rangle(t)$ and $f_{k}(m)$.
+Then, we determine the fully connected feedforward neural network $\mathrm{G}_{\theta}$ by minimizing the loss function $\mathrm{L}(\theta)$:
+$\mathrm{L}(\theta)=\frac{1}{K}\sum_{k=1}^{K}\|\mathrm{G}_{\theta}(\mathbf{I}_{k})-\mathbf{f}_{k}\|^{2},$ (67)
+[…]
+3.2.2 Numerical results
+In our simulations, we let $f_{k}$ be a Gaussian mixture of 3 normal distributions:
+${f_{k}}\propto\sum_{i=1}^{3}w_{i}\mathcal{N}(\mu_{i},\sigma_{i}^{2}),$ (68)
+and sample the parameters $w_{i}$, $\mu_{i}$, and $\sigma_{i}$ as random numbers uniformly on intervals $[0,1]$, $[-0.1,1.1]$, and $[0.05,1.05]$, respectively.
+We solve the mobility-based SIRS model (4) for $\langle I\rangle(t)$ with the initial condition (6) and parameters $\beta=0.8,\varepsilon=0.005$, $\gamma=0.14$, and $I_{0}=0.0001$.
+We sample $K=10,000$ mobility distributions $f_{k}$ in the form of (68) and compute corresponding $\langle I\rangle(t)$.
+We discretize $f_{k}$ and $\langle I\rangle(t)$ with $n_{t}=101$ in (63) and $n_{f}=8$ in (64) and obtain the training set containing $K=10,000$ input-output pairs $(\mathbf{I}_{k},\mathbf{f}_{k})$.
+To improve the performance, we scale the input data $\mathbf{I}_{k}$ by a factor of $50$ and the output data $\mathbf{f}_{k}$ by a factor of $3$.
+The mean values of the input and output sets are around $1$ after rescaling.
+In our simulations, we use a six-layer fully connected neural network, and the width of layers are 128, 512, 512, 512, 128, and 32, respectively.
+Each layer uses a Sigmoid activation function and an $L^{2}$ regularization of $\theta$ with a magnitude of $0.005$.
 For each training step, we sample a random batch from the training set and use the Adam optimizer with a learning rate of 0.0005 to minimize the loss function for each batch.
-We generate a training set of size 10,000 and train the neural network for 2,000 epochs.
-During each training epoch, we optimize the loss function using 1,562 batches with a batch size of 54.
-3.3 Numerical results
-[…]
-Given the dataset $\mathcal{D}$, we obtain the mapping $\mathrm{G}_{\theta}$ by minimizing the loss function (3.2).
-To evaluate performance, we generate an additional six mobility distributions using (62) and simulate the corresponding time series of infected populations $\mathbf{I}$ as inputs.
-Figure 6 shows a comparison between the learned mobility distributions and the ground truth.
-It is noteworthy that the testing inputs $\mathbf{I}_{i}$ are not part of the original training set; however, the inferred mobility distributions are capable of capturing the trends of the ground truth and yield time series that closely match the ground truth.
-(a) Predicted and true mobility distributions $f(m)$
-(b) Predicted and true infected populations $\langle I\rangle(t)$
-Figure 6: (a): A comparison between the predicted mobility distributions and the true distributions. (b): A comparison between the predicted and true $\langle I\rangle(t)$.
-We analyze mobility distributions using real COVID-19 data, shown in Figure 7.
-From February 29th, 2020, to September 16th, 2020, covering the initial 200 days of our dataset, we examine the infection ratio between Massachusetts and New York.
-We calculate the total infected cases by summing daily new infections up to the current time, adjusted with a decaying prefactor to account for previously infected individuals.
-Specifically, we use
-$\langle I\rangle(t)=I_{\text{daily}}(t)+\exp{(-1/\gamma)}\langle I\rangle(t-1).$ (63)
-We recover the daily true case counts $I_{\text{daily}}(t)$ from the death data and fit them with exponential growth for the first 30 days.
-The real data reflects a much smaller scale of disease transmission compared to the training set.
-To use the real data as input, we scale the input by a factor of 50.
-We show the resulting input data in Figure 7(a) and display the learned mobility distributions in Figure 7(b).
+We train the neural network for 2,000 epochs, and during each training epoch, we optimize the loss function using 1,562 batches with a batch size of 54.
+We obtain the mapping $\mathrm{G}_{\theta}$ by minimizing the loss function (67).
+To test the performance of the learned mapping $\mathrm{G}_{\theta}$, we generate 6 additional mobility distributions ($\mathbf{f}_{1},\ldots,\mathbf{f}_{6}$) using (68) and simulate the corresponding time series of infected populations ($\mathbf{I}_{1},\ldots,\mathbf{I}_{6}$) as inputs.
+We compare the predicted outputs $\hat{\mathbf{f}}_{i}=\mathrm{G}_{\theta}(\mathbf{I}_{i})$ for $i=1,\ldots,6$ with the ground truth $\mathbf{f}_{i}$.
+We compute their inverse DCTs and show the comparison between the learned mobility distributions $\hat{f}_{i}$ and the actual distributions ${f}_{i}$ in Figure 6.
+The learned distributions successfully capture the overall trends of the true mobility distributions.
+Figure 6: A comparison of 6 predicted mobility densities $\hat{f}_{i}(m)$ and their ground truth $f_{i}(m)$.
+We solve the mobility-based SIRS model (4) using the predicted mobility $\hat{{f}}_{i}$ and obtain the corresponding time series of the infected population $\langle\hat{I}_{i}\rangle(t)$.
+Figure 7 compares these predicted time series with the ground truth.
+The results show that the inferred mobility distributions generate time series of the infected population that closely match the ground truth.
+Figure 7: A comparison of the predicted infected time series $\hat{\langle I_{i}\rangle}(t)$ and their ground truth $\langle I_{i}\rangle(t)$.
+Next, we focus on inferring mobility distributions using real-world COVID-19 data [15].
+The original dataset includes daily new infection counts for Massachusetts and New York from February 29th to September 16th, 2020.
+We calculate the total number of infected cases by summing the daily new infections with the previously infected cases, adjusted by a decaying factor to account for the recovered cases.
+In particular, we assume that the infected population $\langle I\rangle(t)$ follows
+$\langle I\rangle(t)=I_{\text{daily}}(t)+\exp{(-\gamma)}\langle I\rangle(t-1),$ (69)
+where $I_{daily}(t)$ is the daily infection cases and $\gamma=0.1345$ is the recovery rate, obtained from fitting an SIR model with the Massachusetts data obtained in [15].
+We scale the input $\langle I\rangle(t)$ by a factor of $50$, which is similar to the treatment of the training set.
+We show the rescaled time series of the COVID-19 infected population in Figure 8, together with the inferred mobility distributions from this dataset.
+The real-world data is influenced by numerous factors that have drastically altered social mobility during this period, such as social distancing and lockdown policies.
+Although our model does not account for these factors, the inferred mobility distribution still provides valuable insights.
+We observe that the inferred distributions include a small portion of high-mobility individuals, which accelerates disease transmission during the early stages of the COVID-19 pandemic.
+Classical models, which assume a homogeneous population, often fail to distinguish these high-mobility groups, leading to an overestimation of the pandemic size.
 (a) Real COVID-19 data $\langle I\rangle(t)$
 (b) Predicted mobility $f(m)$
-Figure 7: (a): The scaled ratio of COVID-19 infected cases during the initial wave in Massachusetts and New York from February 29th, 2020, to September 16th, 2020.
-(b): Inferred mobility distributions corresponding to the inputs in (a).
+Figure 8: (a) The COVID-19 infection ratio (scaled by $50$) during the initial wave in Massachusetts and New York from February 29th, 2020, to September 16th, 2020.
+(b) Inferred mobility distributions from the data in (a).
 4 Conclusion and outlook
-In this paper, we integrate social heterogeneity into compartmental models by introducing a mobility variable $m$, assuming it ranges from 00 to $1$ for each individual.
-This approach represents the population in each compartment (susceptible, infected, recovered, etc.) as a density function over $[0,1]$.
-Individuals with higher mobility are more prone to transmitting infectious diseases to others.
-The time evolution of population density for each compartment is described through an integro-differential equation.
-We explore the key properties of the proposed mobility-based compartmental models crucial for epidemiological modeling, such as the basic reproduction number $\mathcal{R}_{0}$, final pandemic size, and mean mobility.
-The proposed mobility-based compartment models also effectively mitigate the overestimation problem inherent in many homogeneous compartmental models when estimating the final pandemic size.
-The basic reproduction rate $\mathcal{R}_{0}$ is often estimated based on the initial growth rate of the infected population.
-In the early stages of a pandemic, individuals with higher mobility tend to get infected quickly, contributing to a higher $\mathcal{R}_{0}$.
-As the pandemic evolves, the recovered population with higher mobility can impart some degree of immunity to the broader population, particularly those with lower mobility.
-This process helps in mitigating the eventual final size of the pandemic.
-We rigorously demonstrate that under the same $\mathcal{R}_{0}$, the classical SIR model yields the largest final pandemic size.
-Numerical findings consistently indicate that a polarized mobility distribution typically results in a significantly smaller final pandemic size compared to the classical SIR model.
-Additionally, we establish a one-to-one correspondence between the time series of the total infected population and the mobility distribution, demonstrating that the mobility distribution can be practically inferred from infection data.
-We propose a deep learning-based algorithm for inferring the mobility distribution directly from the total infected population.
-The primary objective of integrating mobility into compartmental disease models is to utilize the mobility distribution as an intermediary connecting disease dynamics with social dynamics.
-Changes in public health policies and shifts in public opinion about infectious diseases influence disease spread by altering people’s mobility patterns.
-Moving forward, we plan to incorporate opinion dynamics into mobility-based compartmental models to explore the interplay between public opinion and disease transmission dynamics.
+We introduce and analyze density-based compartmental models in epidemiology that incorporate population heterogeneity through mobility.
+We assume that each individual has a mobility variable ranging from 0 to 1 and that individuals with higher mobility are more likely to transmit infectious diseases.
+We describe the disease dynamics by modeling the evolution of mobility distribution functions of the three compartments: susceptible, infected, and recovered.
+Notably, the classical SIR and SIRS models are special cases of our model when the mobility distributions are Dirac delta functions.
+We derive key epidemiological parameters for our mobility-based compartmental models, including the basic reproduction number $\mathcal{R}_{0}$, the final pandemic size, and mean mobility.
+Our models effectively address the overestimation issue common in many homogeneous compartmental models when estimating the final pandemic size.
+Early in a pandemic, individuals with higher mobility tend to become infected more rapidly, resulting in a higher initial growth rate of cases compared to a homogeneous population.
+As the pandemic evolves, the recovered high-mobility population imparts immunity to the broader population, particularly those with lower mobility.
+Most existing models neglect this mechanism, while ours captures it, offering a more accurate estimate of the basic reproduction number and the final pandemic size.
+We prove that the classical SIR model, where the entire population has the same mobility (i.e., mobility distributions are Dirac delta functions), yields the largest final pandemic size compared to models with other mobility distributions.
+Numerical results show that a polarized mobility distribution results in a significantly smaller final pandemic size than the classical SIR model.
+As it is often challenging to measure mobility distributions in practice, we formulate an inference problem to estimate the mobility distribution from the ratio of the total infected population.
+We prove that there is a one-to-one correspondence between the time series of the total infected population and the mobility distribution, which implies that the mobility distribution can be uniquely identified from the infection data.
+Additionally, we propose a numerical approach using a machine learning framework to infer the mobility distribution from the data and validate our results with real COVID-19 data.
+In this paper, we incorporate heterogeneous mobility distributions into the SIR and SIRS models.
+It is also valuable to integrate mobility aspects into more complex compartmental models, such as those with additional compartments or those accounting for birth and death.
+Currently, we assume that each individual’s mobility is time-independent, but one can relax this assumption to consider mobility dynamics influenced by social opinions, public policies, overall infection cases, and many other factors.
+Another potentially interesting extension is to combine the mobility-based compartmental models with control problems and explore how managing population mobility can reduce disease spread.
 Acknowledgement